### v1

# Anomalous dimensions from conformal field theory: generalized $\phi^{2n+1}$ theories

Abstract
We investigate $\phi^{2n+1}$ deformations of the generalized free theory in the $\epsilon$ expansion, where the canonical kinetic term is generalized to a higher-derivative version. For $n=1$, we use the conformal multiplet recombination method to determine the leading anomalous dimensions of the fundamental scalar operator $\phi$ and the bilinear composite operators $\mathcal{J}$. Then we extend the $n=1$ analysis to the Potts model with $S_{N+1}$ symmetry and its higher-derivative generalization, in which $\phi$ is promoted to an $N$-component field. We further examine the Chew-Frautschi plots and their $N$ dependence. However, for each integer $n>1$, the leading anomalous dimensions of $\phi$ and $\mathcal{J}$ are not fully determined and contain one unconstrained constant, which in the canonical cases can be fixed by the results from the traditional diagrammatic method. In all cases, we verify that the multiplet-recombination results are consistent with crossing symmetry using the analytic bootstrap methods.
1 Introduction
For a scalar field theory, the canonical kinetic term is typically proportional to $\partial_{\mu}\phi\partial^{\mu}\phi$, [^1] Summation over repeated indices is implied. with $\phi$ denoting the fundamental scalar. This kinetic term can be rewritten as $\phi\,\Box\phi$ up to a total derivative which we discard. The expression $\phi\,\Box\phi$ can be generalized by employing a higher-order Laplacian. In this paper, we use the word “generalized” to indicate a free theory with a higher-derivative kinetic term and the interacting deformations. The action of the generalized free theory reads
$S\propto\int\mathrm{d}^{d}x\;\phi\,\Box^{k}\phi\,,$ (1)
where $d$ is the spacetime dimension and $k$ is a positive integer (see Brust:2016gjy for more details). One can introduce interaction terms to (1) and investigate the infrared (IR) fixed points of the renormalization group (RG) flow (see for example David:1992vv ; Gracey:2017erc ; Safari:2017irw ; Safari:2017tgs ). We assume that the IR fixed points possess conformal symmetry, allowing for the use of conformal field theory (CFT) methods to study the critical properties.
In a previous paper Guo:2023qtt , we used CFT methods to study the generalized $\phi^{2n}$ theories in the $\epsilon$ expansion. The $\mathbb{Z}_{2}$-odd interactions were not addressed, as $n$ is a positive integer. In this paper, we consider the $\mathbb{Z}_{2}$-odd cases. Namely, we investigate the generalized $\phi^{2n+1}$ theories in the $\epsilon=d_{\text{u}}-d$ expansion, where the upper critical dimensions are given by [^2] It would be interesting to make sense of the corresponding nonperturbative $k>1$ theories, if such theories exist. This is an open problem, to the best our knowledge.
$d_{\text{u}}=2k\,\frac{2n+1}{2n-1}\,.$ (2)
The first instance of these $\mathbb{Z}_{2}$-odd theories corresponds to $n=1$, i.e., the generalized $\phi^{3}$ theory. We also refer to this theory as the generalized Yang-Lee edge singularity, since the canonical $\phi^{3}$ theory describes the Yang-Lee edge singularity Fisher:1978pf .
Before moving on to higher $n$, we also study the generalized $\phi^{3}$ theory with $S_{N+1}$ global symmetry, [^3] One may also introduce an auxiliary field to define an O($N$) symmetric theory in $d=6-\epsilon$ dimensions Fei:2014yja . At large $N$, it is conjectured that the IR CFT at $d=5$ is dual to the Vasiliev higher-spin theory on $\text{AdS}_{6}$. where we promote $\phi$ to an $N$-component field $\vec{\phi}$. The symmetric group $S_{N+1}$ is the group of all permutations of $N+1$ objects. It can also be viewed as the symmetry group of an $N$-simplex. For example, $S_{3}$ is isomorphic to the symmetry group of an equilateral triangle, and this group consists of 6 elements: the identity, rotations by $32\pi/$ 3 and $34\pi/$ 3 about the center, as well as reflections about the three symmetry axes of the equilateral triangle. The mismatch between the number of components of $\phi$ and the degree of the symmetric group originates from the lattice theory described in the next paragraph. In the generalized $S_{N+1}$-symmetric $\phi^{3}$ theory, the $\mathbb{Z}_{2}$ symmetry under the reflection $\vec{\phi}\rightarrow-\vec{\phi}$ is absent, as in the Yang-Lee case. [^4] The symmetric group $S_{N+1}$ contains $\mathbb{Z}_{2}$ subgroups, but the elements of these subgroups do not act by $\vec{\phi}\rightarrow\pm\vec{\phi}$. See appendix C. At $k=1$, this theory describes the $(N+1)$-state Potts model, so we also refer to the generalized $S_{N+1}$-symmetric $\phi^{3}$ theory as the generalized Potts model.
[…]
After discussing the $n=1$ case, we proceed to consider theories with higher $n$ [^9] For studies of the canonical $n=2$ theory with $S_{N+1}$ symmetry, see Codello:2018nbe ; Codello:2020mnt . We only consider the single-field case for the $n>1$ theory in this paper. For $n=2$, the interaction is associated with $\phi^{5}$, and the corresponding theory is related to the Blume-Capel model. More broadly, we study the generalized $\phi^{2n+1}$ theory with an integer $n>1$. A notable feature is that, at leading order in the $\epsilon$ expansion, a family of data is consistent with conformal symmetry, as far as our analysis is concerned. This means that a large set of leading anomalous dimensions is determined up to one parameter, in contrast to the $n=1$ case where the anomalous dimensions are fixed. One motivation for considering the $n>1$ cases is that these theories are associated with multicritical Yang-Lee models vonGehlen:1994rp ; Lencses:2022ira ; Lencses:2023evr ; Lencses:2024wib , and they provide a series of universality classes containing $\mathbb{Z}_{2}$-odd multicritical points. See also Zambelli:2016cbw ; Gracey:2017okb ; Codello:2017qek ; Codello:2017epp for related studies. This concludes our discussion of the theories that will be investigated in this paper.
In these higher-derivative theories, we are interested in the fundamental scalar $\phi$ and the bilinear operators $\mathcal{J}_{\ell}^{(m)}$ of the schematic form $\phi\,\partial^{\ell}\Box^{m}\phi$, which are labeled by the spin $\ell$ and the number of pairs of contracted derivatives $m$. In the free theory, the bilinear operators are (partially) conserved currents, if the spins are sufficiently large. At $m=k-1$, the conserved currents satisfy the usual conservation conditions, i.e., they vanish when contracted with one derivative. For $m<k-1$, the conservation law involves multiple derivatives:
$\overbrace{\partial\ldots\partial}^{c\text{ derivatives}}\!\!\!\!\mathcal{J}_{\ell}^{(m)}=0\,,\qquad\ell\geqslant c\equiv 2(k-m)-1\,,$ (3)
where the indices are suppressed for simplicity. The partially conserved currents do not vanish when we act with fewer derivatives. These operators are related to higher spin symmetry of the free theory. One can construct conserved charges from the partially conserved currents, generalizing the $k=1$ higher spin algebra. We refer to Brust:2016gjy for more details. For $k>1$, the partially conserved currents play a significant role in the analytic bootstrap, as their twists are lower than those of the usual conserved currents. In the lightcone limit, the contribution from the partially conserved currents or their interacting counterparts are more important than the Regge trajectory associated with the conserved currents.
To derive the anomalous dimensions of the bilinear operators, we employ the conformal multiplet recombination method initiated in Rychkov:2015naa , which considered the canonical $\phi^{4}$ theory. In Rychkov:2015naa , it was further suggested that the results can be extended to the canonical $\phi^{3}$ theory at $d=6-\epsilon$. A key difference is the absence of $\mathbb{Z}_{2}$ symmetry in the $\phi^{3}$ theory, in contrast to the $\phi^{4}$ case. Since the procedure in Rychkov:2015naa relies on the $\mathbb{Z}_{2}$ symmetry, its generalization to the $\mathbb{Z}_{2}$-odd cases does not seem straightforward. A similar approach based on the equation of motion was proposed to study the $\phi^{3}$ theory Nii:2016lpa . By applying the equation of motion to correlation functions, the constraints on the anomalous dimensions can be derived more readily. In this paper, we revisit the $\phi^{2n+1}$ theory from the multiplet recombination perspective and extend the results to some spinning operators. As a consistency check, we also verify that the results for the anomalous dimensions and OPE coefficients are consistent with crossing symmetry using the analytic bootstrap.
The multiplet recombination asserts that free multiplets recombine into long multiplets at interacting IR fixed points. For instance, at $d=4-\epsilon$, the Wilson-Fisher multiplet $\{\phi\}$ is a long multiplet into which the free multiplets $\{\phi_{\text{f}}\}$ and $\{\phi^{3}_{\text{f}}\}$ recombine. Here the subscript f implies the free theory. In the Gaussian limit, the descendant $\Box\phi$ is expected to become null and decouple, but a singular change in the normalization can lead to a physical operator, which is identified with the free theory primary operator $\phi^{3}_{\text{f}}$. In other words, the Wilson-Fisher counterpart of $\phi^{3}_{\text{f}}$ is identified as a descendant $\Box\phi$ up to a normalization factor  [^10] All physical operators in the free theory should have counterparts in the interacting theory.:
$\lim_{\epsilon\rightarrow 0}\left(\epsilon^{-1}\Box\phi\right)\propto\phi^{3}_{\text{f}}\,.$ (4)
Despite the resemblance to the equation of motion $\Box\phi\propto\epsilon\,\phi^{3}$, the multiplet recombination is conceptually different, as it can be formulated purely in the CFT language and does not require a Lagrangian description. The multiplet recombination above generalizes to various cases, including the $\mathbb{Z}_{2}$-odd deformations of the free higher-derivative CFTs Gliozzi:2016ysv ; Gliozzi:2017hni ; Gliozzi:2017gzh . See Basu:2015gpa ; Yamaguchi:2016pbj ; Hasegawa:2016piv ; Ghosh:2015opa ; Raju:2015fza ; Gliozzi:2016ysv ; Roumpedakis:2016qcg ; Soderberg:2017oaa ; Behan:2017emf ; Gliozzi:2017hni ; Gliozzi:2017gzh ; Nishioka:2022odm ; Nishioka:2022qmj ; Antunes:2022vtb ; Guo:2023qtt for applications of the multiplet recombination to numerous CFTs, and see Safari:2017irw ; Nii:2016lpa ; Hasegawa:2018yqg ; Skvortsov:2015pea ; Giombi:2016hkj ; Giombi:2017rhm ; Giombi:2016zwa ; Codello:2017qek ; Codello:2018nbe ; Antipin:2019vdg ; Vacca:2019rsh ; Giombi:2020rmc ; Giombi:2020xah ; Dey:2020jlc ; Giombi:2021cnr ; Safari:2021ocb ; Zhou:2022pah ; Bissi:2022bgu ; Herzog:2022jlx ; Giombi:2022vnz ; SoderbergRousu:2023pbe for applications with more emphasis on the equations of motion.
This paper is organized as follows. In section 2, we consider the generalized Yang-Lee edge singularity. We first employ the multiplet recombination to compute the anomalous dimension of $\phi$ and bilinear operators $\mathcal{J}$, then we use the analytic bootstrap to verify that the results are consist with crossing symmetry. In section 3, we carry out a similar analysis, obtaining the anomalous dimensions in the generalized Potts model, and we check that the results are consistent with crossing symmetry as well. Section 4 focuses on the generalized $\phi^{2n+1}$ theory with $n>1$, where we only consider the single-field case. We determine the anomalous dimensions up to a free parameter, then we verify the consistency between the multiplet-recombination results and crossing symmetry. We summarize the main results in table 1. In appendix A, we work out the action of Laplacians on the scalar three-point function. In appendix B, we verify, in all three cases, that the multiplet-recombination results are consistent with crossing symmetry at subleading twist. We briefly review some representation theory of the symmetric group in appendix C. More details about the tensor structures in the Potts model can be found in appendix D. We compute in appendix E the ratios of OPE coefficients relevant to the multiplet recombination. In appendix F, we derive an OPE coefficient used in the analytic bootstrap. Appendix G discusses some special limits in the generalized O($N$) symmetric $\phi^{4}$ theory.
Theories $\gamma_{\phi}$ $\gamma_{\mathcal{J}}$
generalized Yang-Lee edge singularity (22) (29)
generalized Potts model (61) (69)
generalized $\phi^{2n+1}$ theory with $n>1$ (87) (95)
Table 1: Summary of the main results. The second column lists the anomalous dimensions of $\phi$, and the third column lists the anomalous dimensions of the bilinear operators of the schematic form $J\sim\phi\,\partial^{\ell}\Box^{m}\phi$.
2 Generalized Yang-Lee edge singularity
Let us first consider the generalized Yang-Lee CFT. At $d=d_{\text{u}}-\epsilon$, the action reads [^11] We do not need to consider derivative interactions. The reason is that a derivative interaction should be constructed by fewer $\phi$’s, namely two $\phi$’s, but a derivative term with two $\phi$’s is just the kinetic term, up to a total derivative.
$S=\int\mathrm{d}^{d}x\left(\phi\,\Box^{k}\phi+g\mu^{\epsilon/2}\phi^{3}\right)\,,$ (5)
where the upper critical dimension is $d_{\text{u}}=6k$. The parameter $\mu$ has mass dimension 1, so the coupling constant $g$ is dimensionless. We emphasize that $g$ is allowed to be imaginary.
In this section, the anomalous dimensions of $\phi$ and bilinear operators $\mathcal{J}$ are derived from the conformal multiplet recombination. Furthermore, we verify the consistency of the results with crossing symmetry using the analytic bootstrap.
2.1 Multiplet recombination
In the spirit of CFT approaches, we do not refer to the action (5). Instead, the Yang-Lee CFT is characterized by
$\lim_{\epsilon\rightarrow 0}\left(\alpha^{-1}\Box^{k}\phi\right)=\phi^{2}_{\text{f}}\,,$ (6)
where $\alpha=\alpha(\epsilon)$ vanishes in the Gaussian limit: $\lim_{\epsilon\rightarrow 0}\alpha=0$. The above equation means that we identify the descendant $\alpha^{-1}\Box^{k}\phi$ as the deformed version of $\phi^{2}_{\text{f}}$, and the normalization $\alpha^{-1}$ is introduced to ensure a finite Gaussian limit. In other words, the free multiplets $\{\phi_{\text{f}}\}$ and $\{\phi^{2}_{\text{f}}\}$ recombine into a long multiplet $\{\phi\}$.
To study the Yang-Lee CFT, the strategy is to examine the implications of (6) for correlation functions. More specifically, assuming that the $\epsilon\rightarrow 0$ limit is smooth, the condition (6) implies that the Gaussian limit of the correlators in the interacting theory should match the free correlators. Such relations are referred to as matching conditions. Since the deformed CFT data is encoded in the correlation functions, the matching conditions may provide some constraints on the data.
Our focus will be on the two-point and three-point functions. We study the matching conditions for the scalar correlators first. As we will see below, these conditions determine the anomalous dimension of $\phi$ and the normalization $\alpha$, as well as the OPE coefficient associated with $\phi\times\phi\sim\phi$, at leading order in the $\epsilon$ expansion. Then, we consider the matching condition for the spinning correlator involving a primary bilinear operator. [^12] The spin of the bilinear operator can be zero. Using the results for the anomalous dimension of $\phi$ and $\alpha$, we obtain the leading anomalous dimension of the bilinear operator.
The matching condition associated with the two-point function reads
$\lim_{\epsilon\rightarrow 0}\left(\alpha^{-2}\langle\Box^{k}\phi(x_{1})\Box^{k}\phi(x_{2})\rangle\right)=\langle\phi^{2}_{\text{f}}(x_{1})\phi^{2}_{\text{f}}(x_{2})\rangle\,.$ (7)
Namely, the two-point function of the descendant $\alpha^{-1}\Box^{k}\phi$ should reduce to the free correlator of $\phi^{2}_{\text{f}}$ in the Gaussian limit. This leads to the constraint
$\lim_{\epsilon\rightarrow 0}\left[\alpha^{-2}\,4^{2k}\left(\Delta_{\smash{\phi}}\right)_{2k}\left(\Delta_{\smash{\phi}}-\frac{d-2}{2}\right)_{2k}\frac{1}{x_{12}^{2\Delta_{\smash{\phi}}+4k}}\,\right]=\frac{2}{x_{12}^{2\Delta_{\smash{\phi^{2}_{\text{f}}}}}}\,,$ (8)
where $(x)_{y}\equiv\frac{\Gamma(x+y)}{\Gamma(x)}$ is the Pochhammer symbol and $x_{ij}\equiv|x_{i}-x_{j}|$ is the distance between two operators. We have chosen the normalization
$\langle\phi(x_{1})\phi(x_{2})\rangle=1/x_{12}^{\smash{2\Delta_{\phi}}}\,.$ (9)
On the left-hand side of (7), we act on $\langle\phi(x_{1})\phi(x_{2})\rangle$ with the Laplacians, obtaining the two-point function of the descendant $\Box^{k}\phi$. The result follows from the identity
$\Box_{x_{1}}\frac{1}{x_{12}^{2\Delta}}=4\Delta(\Delta-d/2+1)\frac{1}{x_{12}^{2\Delta+2}}\,.$ (10)
On the right-hand side of (7), the free correlator is given by Wick’s theorem. [^13] We assume that the Gaussian limit is smooth, so the normalization in the free theory is also fixed by (9), i.e., $\langle\phi_{\text{f}}(x_{1})\phi_{\text{f}}(x_{2})\rangle=1/x_{12}^{\smash{2\Delta_{\phi_{\text{f}}}}}$. The functional forms in $x_{12}$ indeed match in (8), since
$\lim_{\epsilon\rightarrow 0}\left(2\Delta_{\smash{\phi}}+4k\right)=2\Delta_{\smash{\phi_{\text{f}}}}+4k=8k=2\Delta_{\smash{\phi^{2}_{\text{f}}}}\,.$ (11)
The condition (8) thus imposes a constraint on $\Delta_{\smash{\phi}}$ and $\alpha$. We define the anomalous dimension of $\phi$ as
$\gamma_{\phi}=\Delta_{\smash{\phi}}-\frac{d-2k}{2}\,,$ (12)
and substitute $d=6k-\epsilon$ in (8). Omitting the functional forms in $x_{12}$, we are left with
$(-1)^{k-1}4^{2k}(k-1)!\,k!\left(2k\right)_{2k}\lim_{\epsilon\rightarrow 0}\left(\alpha^{-2}\gamma_{\phi}\right)=2\,,$ (13)
where $\gamma_{\phi}$ arises from the second Pochhammer symbol in (8). For the moment, we note that $\gamma_{\phi}$ is of order $\alpha^{2}$. We will revisit the constraint (13) later.
Now we consider three-point functions. The first matching condition is
$\lim_{\epsilon\rightarrow 0}\left(\alpha^{-1}\langle\Box^{k}\phi(x_{1})\phi(x_{2})\phi(x_{3})\rangle\right)=\langle\phi^{2}_{\text{f}}(x_{1})\phi_{\text{f}}(x_{2})\phi_{\text{f}}(x_{3})\rangle\,.$ (14)
As before, the right-hand side is computed using Wick’s theorem. On the left-hand side, we act on $\langle\phi(x_{1})\phi(x_{2})\phi(x_{3})\rangle$ with the Laplacians. The result is worked out in appendix A. We obtain
$\lim_{\epsilon\rightarrow 0}\left[\alpha^{-1}\sum^{k}_{\begin{subarray}{c}i,j=0\\ i+j\geqslant k\end{subarray}}\left(B_{i,j}\,C_{i,j}^{\Delta_{\smash{\phi}},\Delta_{\smash{\phi}}}\frac{x_{23}^{2(i+j-k)}}{x_{12}^{2i\vphantom{j}}x_{13}^{2j}}\right)\frac{\lambda_{\phi\phi\phi}}{x_{12}^{\Delta_{\smash{\phi}}}x_{13}^{\Delta_{\smash{\phi}}}x_{23}^{\Delta_{\smash{\phi}}}}\right]=\frac{2}{x_{12}^{\Delta_{\smash{\phi^{2}_{\text{f}}}}}x_{13}^{\Delta_{\smash{\phi^{2}_{\text{f}}}}}}\,,$ (15)
where $\lambda_{\phi\phi\phi}$ is the OPE coefficient of $\phi\in\phi\times\phi$. The coefficient consists of a part independent of the scaling dimensions:
$B_{i,j}$ $=\frac{(-1)^{i+j-k}2^{2k}k!}{(k-i)!\,(k-j)!\,(i+j-k)!}\,,$ (16)
and another part given by
$C_{i,j}^{a,b}$ $=\Big{(}\frac{a}{2}\Big{)}_{i}\Big{(}\frac{b}{2}\Big{)}_{j}\left(\frac{a+b}{2}-\frac{d-2}{2}+i+j-k\right)_{2k-i-j}\,.$ (17)
Due to the last Pochhammer symbol, $C_{i,j}^{\Delta_{\smash{\phi}},\Delta_{\smash{\phi}}}$ is proportional to the anomalous dimension $\gamma_{\phi}$, except for the $i=j=k$ case where $C_{i,j}^{\Delta_{\smash{\phi}},\Delta_{\smash{\phi}}}$ is finite in the Gaussian limit. In other words, the dominant term in the $\epsilon\rightarrow 0$ limit is given by $i=j=k$. Then, one can check that the functional forms match in (15). We obtain a constraint on $\lambda_{\phi\phi\phi}$ and $\alpha$:
$(-1)^{k}2^{2k}\left(k\right)_{k}^{2}\lim_{\epsilon\rightarrow 0}\left(\alpha^{-1}\lambda_{\phi\phi\phi}\right)=2\,,$ (18)
indicating that the OPE coefficient $\lambda_{\phi\phi\phi}$ is of order $\alpha$. Another matching condition is
$\lim_{\epsilon\rightarrow 0}\left(\alpha^{-3}\langle\Box^{k}\phi(x_{1})\Box^{k}\phi(x_{2})\Box^{k}\phi(x_{3})\rangle\right)=\langle\phi^{2}_{\text{f}}(x_{1})\phi^{2}_{\text{f}}(x_{2})\phi^{2}_{\text{f}}(x_{3})\rangle\,.$ (19)
The action of the Laplacians is derived in appendix A. We have
$\lim_{\epsilon\rightarrow 0}\Bigg{\{}\alpha^{-3}$ $\sum^{k}_{\begin{subarray}{c}i_{1},j_{1}=0\\ i_{1}+j_{1}\geqslant k\end{subarray}}\;\;\sum^{k}_{\begin{subarray}{c}i_{2},j_{2}=0\\ i_{2}+j_{2}\geqslant k\end{subarray}}\;\;\sum^{k}_{\begin{subarray}{c}i_{3},j_{3}=0\\ i_{3}+j_{3}\geqslant k\end{subarray}}\Bigg{[}B_{i_{1},j_{1}}C_{i_{1},j_{1}}^{\Delta_{\smash{\phi}},\Delta_{\smash{\phi}}}B_{i_{2},j_{2}}C_{i_{2},j_{2}}^{\Delta_{\smash{\phi}}+2i_{1},\Delta_{\smash{\phi}}+2(k-i_{1}-j_{1})}$
×$\times B_{i_{3},j_{3}}C_{i_{3},j_{3}}^{\Delta_{\smash{\phi}}+2(j_{2}+k-i_{1}-j_{1}),\Delta_{\smash{\phi}}+2(j_{1}+k-i_{2}-j_{2})}x_{12}^{2(i_{3}+j_{3}-i_{1}-i_{2}-k)}$
×$\times x_{13}^{2(i_{2}+j_{2}-j_{1}-j_{3}-k)}x_{23}^{2(i_{1}+j_{1}-j_{2}-i_{3}-k)}\Bigg{]}\frac{\lambda_{\phi\phi\phi}}{x_{12}^{\Delta_{\smash{\phi}}}x_{13}^{\Delta_{\smash{\phi}}}x_{23}^{\Delta_{\smash{\phi}}}}\Bigg{\}}=\frac{8}{x_{12}^{\Delta_{\smash{\phi^{2}_{\text{f}}}}}x_{13}^{\Delta_{\smash{\phi^{2}_{\text{f}}}}}x_{23}^{\Delta_{\smash{\phi^{2}_{\text{f}}}}}}\,,$ (20)
where $B_{i,j}$ and $C_{i,j}^{a,b}$ are defined in (16) and (17). As before, the functional forms match, leaving us with a constraint on $\gamma_{\phi}$, $\lambda_{\phi\phi\phi}$ and $\alpha$:
$(-1)^{k-1}2^{12k-5}(3k-1)!\left(1/2\right)_{k}^{3}\lim_{\epsilon\rightarrow 0}\left[\alpha^{-3}(\epsilon-6\gamma_{\phi})\lambda_{\phi\phi\phi}\right]=8\,.$ (21)
If we divide the left-hand side by those of (13) and (18), then the right-hand side becomes 2, so $\lim_{\epsilon\rightarrow 0}(\epsilon\,\gamma_{\phi}^{-1})$ is finite for positive integer $k$ and $\gamma_{\phi}$ is of order $\epsilon$. It follows that $\alpha$ and $\lambda$ are of order $\epsilon^{1/2}$. Moreover, we derive the explicit expression for the leading anomalous dimension
$\gamma_{\phi}=\frac{\epsilon}{2\left(3+\frac{(-1)^{k}k!\left(3k\right)_{k}}{2^{4k-4}\left(1/2\right)_{k}^{2}}\right)}+O(\epsilon^{3/2})\,,$ (22)
which generalizes the $k=1$ result in Fisher:1978pf . For $k=1$, the anomalous dimension of $\phi$ is negative: $\gamma_{\phi}=-\epsilon/18$, implying that the theory is non-unitary.  [^14] For $k>1$, the generalized free theory itself is non-unitary due to the violation of the unitarity bound. We also obtain
$\alpha^{2}$ $=-\frac{(2k)!\,(3k-1)!\,\epsilon}{2^{5-4k}k!+6(-1)^{k}\frac{\left(1/2\right)_{k}^{2}}{\left(3k\right)_{k}}}+O(\epsilon^{3/2})\,,$ (23)
$\lambda_{\phi\phi\phi}$ $=\frac{(-1)^{k}\alpha}{2^{6k-5}(3/2)_{k-1}^{2}}+O(\epsilon)\,.$ (24)
At leading order, the OPE coefficient $\lambda_{\phi\phi\phi}$ and $\alpha$ are imaginary if $k$ is even or $k\in\{1,3\}$.
Now we consider the spinning correlator involving a generic primary bilinear operator $\mathcal{J}_{\ell}^{(m)}\sim\phi\partial^{\ell}\Box^{m}\phi$ [^15] The operators $\mathcal{J}_{\ell}^{(m)}$ should be nondegenerate. See, for example, footnote 12 of Guo:2023qtt . The matching condition is
$\lim_{\epsilon\rightarrow 0}\left(\alpha^{-2}\langle\Box^{k}\phi(x_{1})\Box^{k}\phi(x_{2})\mathcal{J}_{\ell}^{(m)}(x_{3},z)\rangle\right)=\langle\phi^{2}_{\text{f}}(x_{1})\phi^{2}_{\text{f}}(x_{2})\mathcal{J}^{(m)}_{\ell,\text{f}}(x_{3},z)\rangle\,.$ (25)
We have used the index-free notation where the spin indices are contracted with those of the polarization vector $z\in\mathbb{C}^{d}$. We simplify the calculation further by focusing on the leading term in the $x_{3}\rightarrow\infty$ expansion Gliozzi:2017gzh . This makes it easier to extract the constraint given by (25), since the action of the Laplacians on the leading term is simpler. See section 3 of Guo:2023qtt for more details on the treatment of the spinning correlator. We have
$\lim_{\epsilon\rightarrow 0}\Bigg{[}$ $\alpha^{-2}4^{2k}\left(\Delta_{\smash{\phi}}-\frac{\Delta_{\mathcal{J}}}{2}+\frac{\ell}{2}\right)_{2k}\left(\Delta_{\smash{\phi}}-\frac{\Delta_{\mathcal{J}}}{2}-\frac{\ell}{2}-\frac{d-2}{2}\right)_{2k}$
×$\times\frac{\lambda_{\phi\phi\mathcal{J}}\left(z\cdot x_{12}\right)^{\ell}}{x_{12}^{2\Delta_{\smash{\phi}}+4k-\Delta_{\mathcal{J}}+\ell}|x_{3}|^{2\Delta_{\mathcal{J}}}}+O(\,|x_{3}|^{-2\Delta_{\mathcal{J}}-1})\Bigg{]}=\frac{\lambda_{\phi^{2}\phi^{2}\mathcal{J},\text{f}}\left(z\cdot x_{12}\right)^{\ell}}{x_{12}^{2\Delta_{\smash{\phi^{2}_{\text{f}}}}-\Delta_{\smash{\mathcal{J}_{\text{f}}}}+\ell}|x_{3}|^{2\Delta_{\smash{\mathcal{J}_{\text{f}}}}}}+O(\,|x_{3}|^{-2\Delta_{\smash{\mathcal{J}_{\text{f}}}}-1})\,,$ (26)
where $\lambda_{\phi\phi\mathcal{J}}$ and $\lambda_{\phi^{2}\phi^{2}\mathcal{J},\text{f}}$ are the OPE coefficients of $\mathcal{J}\in\phi\times\phi$ and $\mathcal{J}_{\text{f}}\in\phi^{2}_{\text{f}}\times\phi^{2}_{\text{f}}$. The anomalous dimensions of bilinear operators are defined by
$\gamma_{\mathcal{J}}=\Delta_{\mathcal{J}}-(d-2k+2m+\ell)\,,$ (27)
where the terms in parentheses correspond to the classical dimension of $\mathcal{J}_{\ell}^{(m)}$. The functional forms in (2.1) match. We then write
$4^{2k}(2k-m-1)!\,(-m)_{m}(1-3k-\ell-m)_{2k}\lim_{\epsilon\rightarrow 0}\left[\alpha^{-2}\left(\gamma_{\phi}-\frac{\gamma_{\mathcal{J}}}{2}\right)\right]=\frac{\lambda_{\phi^{2}\phi^{2}\mathcal{J},\text{f}}}{\lambda_{\phi\phi\mathcal{J},\text{f}}}\,,$ (28)
where $\lim_{\epsilon\rightarrow 0}\lambda_{\phi\phi\mathcal{J}}=\lambda_{\phi\phi\mathcal{J},\text{f}}$ is used. Using the ratio of OPE coefficients $\frac{\lambda_{\phi^{2}\phi^{2}\mathcal{J},\text{f}}}{\lambda_{\phi\phi\mathcal{J},\text{f}}}=4$ derived in appendix E, we obtain
$\gamma_{\mathcal{J}}=2\gamma_{\phi}-\frac{\alpha^{2}}{2^{4k-3}(2k-m-1)!(-m)_{m}(1-3k-\ell-m)_{2k}}+O(\epsilon^{3/2})\,,$ (29)
where $\gamma_{\phi}$ and $\alpha^{2}$ are given in (22) and (23). This is the $k\geqslant 1$ generalization of the $k=1$ result in Giombi:2016hkj . The stress tensor is the spin-2 current on the highest trajectory $m=k-1$, and its anomalous dimension vanishes to order $\epsilon$, according to (29).
The spin-0 operator on the lowest trajectory $\mathcal{J}_{\ell=0}^{(m=0)}=\phi^{2}$ is not a primary, so the matching condition (25) does not apply in this case. Nevertheless, we can set $\ell=m=0$ in (29). It turns out that we obtain the shadow of $\phi$:
$\Delta_{\smash{\phi}}+\Delta_{\mathcal{J}}\big{|}_{\ell=m=0}=d+O(\epsilon^{3/2})\,,$ (30)
which holds for $k\geqslant 1$. The $k=1$ case was already noticed in Goncalves:2018nlv .
2.2 Analytic bootstrap
In this subsection, we verify the consistency between the results above and crossing symmetry. In the lightcone limit of the crossing equation for the four-point function $\langle\phi\phi\phi\phi\rangle$, the anomalous dimensions of the bilinear operators are read off from the terms with enhanced singularity in the crossed channel. [^16] The meaning of the enhanced singularity is as follows. An expression with an enhanced singularity is power-law divergent after repeated action of the conformal Casimir operator Alday:2015eya . The results are consistent with those derived from the multiplet recombination.
The four-point function of identical scalars takes the form
$\langle\phi(x_{1})\phi(x_{2})\phi(x_{3})\phi(x_{4})\rangle=\frac{g(z,\bar{z})}{x_{12}^{2\Delta_{\smash{\phi}}}x_{34}^{2\Delta_{\smash{\phi}}}}\,,$ (31)
where $z$ and $\bar{z}$ are defined by
$z\bar{z}=\frac{x^{2}_{12}x^{2}_{34}}{x^{2}_{13}x^{2}_{24}}\,,\qquad(1-z)(1-\bar{z})=\frac{x^{2}_{14}x^{2}_{23}}{x^{2}_{13}x^{2}_{24}}\,.$ (32)
The crossing equation is
$g(z,\bar{z})=\frac{z^{\Delta_{\smash{\phi}}}\bar{z}^{\Delta_{\smash{\phi}}}}{(1-z)^{\Delta_{\smash{\phi}}}(1-\bar{z})^{\Delta_{\smash{\phi}}}}g(1-\bar{z},1-z)\,.$ (33)
The conformal block decomposition of the left-hand side reads
$g(z,\bar{z})=\sum_{\mathcal{O}}P_{\mathcal{O}}G_{\mathcal{O}}(z,\bar{z})\,,$ (34)
where $P_{\mathcal{O}}$ is the squared OPE coefficient of $\mathcal{O}$ and $G_{\mathcal{O}}$ is the conformal block associated with $\mathcal{O}$. We consider the lightcone limit $z\rightarrow 0$. The leading term of the conformal block in the lightcone expansion is
$G_{\mathcal{O}}(z,\bar{z})=z^{(\Delta_{\mathcal{O}}-\ell)/2}k_{h}(\bar{z})+O(z^{(\Delta_{\mathcal{O}}-\ell)/2+1})\,.$ (35)
Here $h=\frac{\Delta_{\mathcal{O}}+\ell}{2}$ is the conformal spin and $k_{h}(\bar{z})$ denotes the $\mathbb{R}$) block
$k_{h}(\bar{z})=\bar{z}^{h}{}_{2}F_{1}(h,h;2h;\bar{z})\,,$ (36)
where ${}_{2}F_{1}(a,b;c;\bar{z})$ is the Gaussian hypergeometric function. In this convention, the OPE coefficients have a different normalization from those in the multiplet recombination approach. They are related by two-point function coefficients $\lambda_{\mathcal{O}\mathcal{O}\mathds{1}}$:
$P_{\mathcal{O}}=\frac{\lambda_{\phi\phi\mathcal{O}}^{2}}{\lambda_{\mathcal{O}\mathcal{O}\mathds{1}}}\,.$ (37)
The right-hand side of (33) has the conformal block decomposition
$g(1-\bar{z},1-z)=1+P_{\phi}G_{\phi}(1-\bar{z},1-z)+\ldots\,,$ (38)
where 1 is the contribution of the identity operator and the second term is the contribution of $\phi$. The ellipsis represents terms of higher order in $\epsilon$, and they will be omitted below. At the leading order in the $z\rightarrow 0$ expansion, the right-hand side of (33) implies an enhanced singularity in the limit $\bar{z}\rightarrow 1$. This behavior matches the infinite sums of the $\mathbb{R}$) blocks on the left-hand side of (33) with $\mathcal{O}=\mathcal{J}^{(m=0)}_{\ell}$, thanks to the asymptotic relation Simmons-Duffin:2016wlq
$\sum_{\ell=0,2,4,\ldots}2S_{p}(h)\,k_{h}(\bar{z})\sim\left(\frac{1-\bar{z}}{\bar{z}}\right)^{p}\,,\qquad(\bar{z}\rightarrow 1)\,,$ (39)
where the coefficient is
$S_{p}(h)=\frac{\Gamma(h-p-1)}{\Gamma(-p)^{2}\Gamma(h+p+1)}\frac{\Gamma(h)^{2}}{\Gamma(2h-1)}\,.$ (40)
The analysis can be extended to higher-twist bilinear operators, i.e., to the case of $m>0$. In appendix B, we also verify the multiplet-recombination results at $m=1$ using the analytic bootstrap.
We expand the left-hand side of (33) in the anomalous dimension
$\tilde{\gamma}_{\mathcal{J}}=\Delta_{\mathcal{J}}-\left(2\Delta_{\smash{\phi}}+2m+\ell\right)\,,$ (41)
where $\Delta_{\smash{\phi}}$ is the full scaling dimension of $\phi$. The conformal block associated with $\mathcal{J}_{\ell}^{(m=0)}$ becomes
$G_{\mathcal{J}}(z,\bar{z})=z^{\Delta_{\smash{\phi}}}\left(k_{h}(\bar{z})+\frac{1}{2}\tilde{\gamma}_{\mathcal{J}}k_{h}(\bar{z})\log z+O(\tilde{\gamma}_{\mathcal{J}}^{2})\right)+O(z^{(\Delta_{\mathcal{J}}-\ell)/2+1})\,,$ (42)
where we have suppressed the indices associated with $\mathcal{J}_{\ell}^{(0)}$ for simplicity. Consider the right-hand side of the crossing equation (33) at leading order in the lightcone expansion. The identity operator contributes to the enhanced singularity at order $\epsilon^{0}$. The corresponding term is $(\frac{1-\bar{z}}{\bar{z}})^{-\Delta_{\smash{\phi_{\text{f}}}}}$, with $\Delta_{\smash{\phi_{\text{f}}}}=2k$. The free squared OPE coefficients of $\mathcal{J}^{(m=0)}_{\ell}$ can be derived using (39):
$P_{\smash{\mathcal{J}^{(0)}_{\ell}},\text{f}}=2S_{-2k}(h)=2\frac{(2k)_{\ell}^{2}}{\ell!\,(4k+\ell-1)_{\ell}}\,.$ (43)
The leading correction on the right-hand side of (33) is of order $\epsilon^{1}$, as $P_{\phi}\sim\epsilon^{1}$ according to (24) and (37). So we conclude that $\tilde{\gamma}_{\mathcal{J}}\sim\epsilon^{1}$ in (42), which is necessary for the consistency of (33). At order $\epsilon^{1}$, the enhanced singularity is associated with the conformal block $G_{\phi}(1-\bar{z},1-z)$ in the crossed channel. To extract the anomalous dimensions $\tilde{\gamma}_{\mathcal{J}}$ in (42), we consider the $\log z$ part of a scalar block Li:2020ijq
$G_{\mathcal{O}}(1-\bar{z},1-z)\Big{|}_{\log z\text{ part}}\!\!=-\frac{\Gamma(\Delta_{\smash{\mathcal{O}}})}{\Gamma(\Delta_{\smash{\mathcal{O}}}/2)^{2}}(1-\bar{z})^{\frac{\Delta_{\smash{\mathcal{O}}}}{2}}{}_{2}F_{1}\!\left(\frac{\Delta_{\smash{\mathcal{O}}}}{2},\frac{\Delta_{\smash{\mathcal{O}}}}{2};\Delta_{\smash{\mathcal{O}}}-\frac{d-2}{2};1-\bar{z}\right)+O(z)\,,$ (44)
where $\mathcal{O}$ is a scalar operator. We focus only on the terms with the enhanced singularity at order $\epsilon^{1}$:
$\frac{\bar{z}^{\Delta_{\smash{\phi}}}}{(1-\bar{z})^{\Delta_{\smash{\phi}}}}P_{\phi}G_{\phi}(1-\bar{z},1-z)\Big{|}_{\log z\text{ part}}$
[…]
where the Casimir-regular terms do not contribute to the enhanced singularity. Using (24), (37), (39) and (43), we obtain the anomalous dimensions $\tilde{\gamma}_{\mathcal{J}}$. The result agrees with (29), confirming the consistency of the multiplet-recombination results (24) and (29) with crossing symmetry.
Due to the argument $\Delta_{\smash{\mathcal{O}}}-\frac{d-2}{2}$, the hypergeometric function in (44) with $\mathcal{O}=\phi$ diverges as $1/\gamma_{\phi}$ in the Gaussian limit. Nonetheless, the conformal block decomposition (38) makes sense in the $\epsilon$ expansion, because $P_{\phi}/\gamma_{\phi}$ is finite as $\epsilon\rightarrow 0$. Using (22), we find that the finite term is exactly equal to the contribution of $\phi^{2}_{\text{f}}$ in the free theory:
$\lim_{\epsilon\rightarrow 0}\left[P_{\phi}G_{\phi}(1-\bar{z},1-z)\Big{|}_{\log z\text{ part}}\,\right]=P_{\phi^{2}_{\text{f}}}\,G_{\phi^{2}_{\text{f}}}(1-\bar{z},1-z)\Big{|}_{\log z\text{ part}}\,,$ (46)
which is compatible with the multiplet recombination.
3 Generalized Potts model
We can extend the $\phi^{3}$ discussion to the case of an $N$-component field $\phi_{a}$, where $a=1,2,\ldots,N$. The action that describes the ($N+1$)-state Potts model is
$S=\int\mathrm{d}^{d}x\left(\sum_{a=1}^{N}\phi_{a}\Box^{k}\phi_{a}+g\mu^{\epsilon/2}\sum_{a,b,c=1}^{N}T_{abc}\,\phi_{a}\phi_{b}\phi_{c}\right)\,,$ (47)
where $T_{abc}$ is an $S_{N+1}$ invariant tensor [^17] The rank-3 invariant tensor of $S_{N+1}$ is unique up to a scalar multiple wallace1978spin .
$T_{abc}=\sum_{\alpha=1}^{N+1}e_{a}^{\alpha}e_{b}^{\alpha}e_{c}^{\alpha}\,.$ (48)
As mentioned in the introduction, the vectors $e_{a}^{1},e_{a}^{2}\ldots,e_{a}^{N+1}$ have the same magnitude and any two distinct vectors $e_{a}^{\alpha}$ and $e_{a}^{\beta}$ form the same angle. In fact, these vectors are associated with the vertices of an $N$-simplex, whose symmetry group is $S_{N+1}$. The field $\phi_{a}$ lives in an $N$-dimensional irreducible representation of $S_{N+1}$ called the standard representation. More details can be found in appendix C.
In this section, we obtain the anomalous dimension of $\phi_{a}$ and bilinear operators $\mathcal{J}_{\ell}^{(m)}$, using the multiplet recombination. For positive integers $N>2$, there are four types of bilinear operators, due to the tensor product decomposition murnaghan1938analysis
$(N,1)\otimes(N,1)=(N+1)+(N,1)+(N-1,2)+(N-1,1^{2})\,.$ (49)
Here the partitions of $N+1$ boxes represent the Young diagrams:
We refer to these four representations as the trivial, standard, symmetric traceless and antisymmetric representations (see appendix C for an explanation of these names). The general results for the anomalous dimensions depend on $N$, and some particularly interesting cases are the percolation theory ($N\rightarrow 0$) and the spanning forest ($N\rightarrow-1$). As in subsection 2.2, we verify the recombination results using the analytic bootstrap.
3.1 Multiplet recombination
The multi-field version of (6) takes the form
$\lim_{\epsilon\rightarrow 0}\left(\alpha^{-1}\Box^{k}\phi_{a}\right)=\sum_{b,c=1}^{N}T_{abc}\,\phi_{b,\text{f}}\,\phi_{c,\text{f}}\,,$ (50)
where $T_{abc}$ is given in (48). Below, we derive the anomalous dimensions of $\phi_{a}$ and the four types of bilinear operators using (50). The procedure is similar to that of section 2.1. The difference is that here the correlators have tensor structures related to the $S_{N+1}$ symmetry, and the constraint equations are modified by $N$-dependent factors.
The matching condition analogous to (7) reads
$\lim_{\epsilon\rightarrow 0}\left(\alpha^{-2}\langle\Box^{k}\phi_{a_{1}}(x_{1})\Box^{k}\phi_{a_{2}}(x_{2})\rangle\right)=\langle K_{a_{1},\text{f}}(x_{1})K_{a_{2},\text{f}}(x_{2})\rangle\,,$ (51)
where we have defined
$K_{a,\text{f}}\equiv\sum_{b,c=1}^{N}T_{abc}\,\phi_{b,\text{f}}\,\phi_{c,\text{f}}\,.$ (52)
The left-hand side of (51) is nearly identical to that of (8). The difference is that here we have a tensor structure $\delta_{a_{1}a_{2}}$, which is the only rank-2 $S_{N+1}$-invariant tensor (up to a scalar multiple). [^18] The uniqueness can be seen from the tensor product decomposition (49), where the trivial representation occurs only once on the right-hand side wallace1978spin . The correlator on the right-hand side is again given by Wick’s theorem. We obtain
$\lim_{\epsilon\rightarrow 0}\left[\alpha^{-2}\,4^{2k}\left(\Delta_{\smash{\phi}}\right)_{2k}\left(\Delta_{\smash{\phi}}-\frac{d-2}{2}\right)_{2k}\frac{\delta_{a_{1}a_{2}}}{x_{12}^{2\Delta_{\smash{\phi}}+4k}}\,\right]=\frac{2(N-1)(N+1)^{2}\,\delta_{a_{1}a_{2}}}{x_{12}^{2\Delta_{\smash{K_{\text{f}}}}}}\,,$ (53)
where $\Delta_{K_{\text{f}}}$ denotes the scaling dimension of $K_{a,\text{f}}$. We have used (126) and (48), which lead to $\delta_{a_{1}a_{2}}$ and the $N$-dependent coefficient on the right-hand side. The functional forms and the tensor structures match, so we are left with the constraint
$(-1)^{k-1}4^{2k}(k-1)!\,k!\left(2k\right)_{2k}\lim_{\epsilon\rightarrow 0}\left(\alpha^{-2}\gamma_{\phi}\right)=2(N-1)(N+1)^{2}\,,$ (54)
where $\gamma_{\phi}$ is the anomalous dimension of $\phi_{a}$ defined by (12).
For three-point functions, we consider the matching condition
$\lim_{\epsilon\rightarrow 0}\left(\alpha^{-1}\langle\Box^{k}\phi_{a_{1}}(x_{1})\phi_{a_{2}}(x_{2})\phi_{a_{3}}(x_{3})\rangle\right)=\langle K_{a_{1},\text{f}}(x_{1})\,\phi_{a_{2},\text{f}}(x_{2})\phi_{a_{3},\text{f}}(x_{3})\rangle\,.$ (55)
On the left-hand side, the tensor structure is given by $T_{a_{1}a_{2}a_{3}}$, which is the only rank-3 $S_{N+1}$-invariant tensor (up to a scalar multiple). On the right-hand side, the free correlator also has the tensor structure $T_{a_{1}a_{2}a_{3}}$. We obtain
$\lim_{\epsilon\rightarrow 0}\left[\alpha^{-1}\sum^{k}_{\begin{subarray}{c}i,j=0\\ i+j\geqslant k\end{subarray}}\left(B_{i,j}\,C_{i,j}^{\Delta_{\smash{\phi}},\Delta_{\smash{\phi}}}\frac{x_{23}^{2(i+j-k)}}{x_{12}^{2i\vphantom{j}}x_{13}^{2j}}\right)\frac{\lambda_{\phi\phi\phi}T_{a_{1}a_{2}a_{3}}}{x_{12}^{\Delta_{\smash{\phi}}}x_{13}^{\Delta_{\smash{\phi}}}x_{23}^{\Delta_{\smash{\phi}}}}\right]=\frac{2\,T_{a_{1}a_{2}a_{3}}}{x_{12}^{\Delta_{\smash{K_{\text{f}}}}}x_{13}^{\Delta_{\smash{K_{\text{f}}}}}}\,,$ (56)
where $B_{i,j}$ and $C_{i,j}$ are the same as in (16) and (17). The constraint on $\lambda_{\phi\phi\phi}$ and $\alpha$ takes the same form as (18):
$(-1)^{k}2^{2k}(k)_{k}^{2}\lim_{\epsilon\rightarrow 0}\left(\alpha^{-1}\lambda_{\phi\phi\phi}\right)=2\,.$ (57)
Next, we consider the matching condition
$\lim_{\epsilon\rightarrow 0}\left(\alpha^{-3}\langle\Box^{k}\phi_{a_{1}}(x_{1})\Box^{k}\phi_{a_{2}}(x_{2})\Box^{k}\phi_{a_{3}}(x_{3})\rangle\right)=\langle K_{a_{1},\text{f}}(x_{1})K_{a_{2},\text{f}}(x_{2})K_{a_{3},\text{f}}(x_{3})\rangle\,,$ (58)
which results in
$\lim_{\epsilon\rightarrow 0}\Bigg{\{}\alpha^{-3}$ $\sum^{k}_{\begin{subarray}{c}i_{1},j_{1}=0\\ i_{1}+j_{1}\geqslant k\end{subarray}}\;\;\sum^{k}_{\begin{subarray}{c}i_{2},j_{2}=0\\ i_{2}+j_{2}\geqslant k\end{subarray}}\;\;\sum^{k}_{\begin{subarray}{c}i_{3},j_{3}=0\\ i_{3}+j_{3}\geqslant k\end{subarray}}\Bigg{[}B_{i_{1},j_{1}}C_{i_{1},j_{1}}^{\Delta_{\smash{\phi}},\Delta_{\smash{\phi}}}B_{i_{2},j_{2}}C_{i_{2},j_{2}}^{\Delta_{\smash{\phi}}+2i_{1},\Delta_{\smash{\phi}}+2(k-i_{1}-j_{1})}$
×$\!\!\times B_{i_{3},j_{3}}C_{i_{3},j_{3}}^{\Delta_{\smash{\phi}}+2(j_{2}+k-i_{1}-j_{1}),\Delta_{\smash{\phi}}+2(j_{1}+k-i_{2}-j_{2})}x_{12}^{2(i_{3}+j_{3}-i_{1}-i_{2}-k)}$
×$\!\!\times x_{13}^{2(i_{2}+j_{2}-j_{1}-j_{3}-k)}x_{23}^{2(i_{1}+j_{1}-j_{2}-i_{3}-k)}\Bigg{]}\frac{\lambda_{\phi\phi\phi}T_{a_{1}a_{2}a_{3}}}{x_{12}^{\Delta_{\smash{\phi}}}x_{13}^{\Delta_{\smash{\phi}}}x_{23}^{\Delta_{\smash{\phi}}}}\Bigg{\}}=\frac{8(N-2)(N+1)^{2}\,T_{a_{1}a_{2}a_{3}}}{x_{12}^{\Delta_{\smash{K_{\text{f}}}}}x_{13}^{\Delta_{\smash{K_{\text{f}}}}}x_{23}^{\Delta_{\smash{K_{\text{f}}}}}}\,.$ (59)
In the $\epsilon$ expansion, this leads to the constraint
$(-1)^{k-1}2^{12k-5}(3k-1)!\left(1/2\right)_{k}^{3}\lim_{\epsilon\rightarrow 0}\left[\alpha^{-3}(\epsilon-6\gamma_{\phi})\lambda_{\phi\phi\phi}\right]=8(N-2)(N+1)^{2}\,.$ (60)
Combining (54), (57) and (60), we obtain the anomalous dimension
$\gamma_{\phi}=\frac{\epsilon}{2\left(3+\frac{N-2}{N-1}\frac{(-1)^{k}k!(3k)_{k}}{2^{4k-4}(1/2)_{k}^{2}}\right)}+O(\epsilon^{3/2})\,.$ (61)
This expression is very similar to (22), but here an $N$-dependent factor appears in the second term in the denominator. [^19] Interestingly, the anomalous dimension of $\phi$ is $k$-independent to order $\epsilon$ at $N=2$: $\gamma_{\phi}=\epsilon/6+O(\epsilon^{3/2})$. At $k=1$, the result agrees with that from the RG analysis Amit:1976pz . We also obtain $\alpha$ and the OPE coefficient of $\phi$:
$\alpha^{2}$ $=-\frac{(2k)!\,(3k-1)!\,\epsilon}{(N+1)^{2}\left((N-2)2^{5-4k}k!+6(N-1)(-1)^{k}\frac{(1/2)_{k}^{2}}{(3k)_{k}}\right)}+O(\epsilon^{3/2})\,,$ (62)
$\lambda_{\phi\phi\phi}$ $=\frac{(-1)^{k}\alpha}{2^{6k-5}(3/2)_{k-1}^{2}}+O(\epsilon^{3/2})\,.$ (63)
As mentioned in the introduction, the $N\rightarrow\infty$ limit yields the decoupled Yang-Lee theory. One can check that
$\lim_{N\rightarrow\infty}\gamma_{\phi}^{\text{Potts}}=\gamma_{\phi}^{\text{YL}}\,,$ (64)
where $\gamma_{\phi}^{\text{Potts}}$ and $\gamma_{\phi}^{\text{YL}}$ are given in (61) and (22). Moreover, we can take the $N\rightarrow 0,-1$ limits of (61) to study the generalized percolation theory and spanning forest. The anomalous dimensions are
$\gamma_{\phi}\big{|}_{N\rightarrow 0}=\frac{\epsilon}{2\left(3+\frac{(-1)^{k}k!(3k)_{k}}{2^{4k-5}(1/2)_{k}^{2}}\right)}+O(\epsilon^{3/2})\,,\qquad\gamma_{\phi}\big{|}_{N\rightarrow-1}=\frac{\epsilon}{2\left(3+\frac{3(-1)^{k}k!(3k)_{k}}{2^{4k-3}(1/2)_{k}^{2}}\right)}+O(\epsilon^{3/2})\,.$ (65)
At $k=1$, they lead to the critical exponents of percolation theory $\eta_{\text{percolation}}=-\epsilon/21+O(\epsilon^{3/2})$ and spanning forest $\eta_{\text{spanning forest}}=-\epsilon/15+O(\epsilon^{3/2})$.
Now, we consider the bilinear operators. The matching condition reads
$\lim_{\epsilon\rightarrow 0}\left(\alpha^{-2}\langle\Box^{k}\phi_{a_{1}}(x_{1})\Box^{k}\phi_{a_{2}}(x_{2})\mathcal{J}_{\ell}^{(m)}(x_{3},z)\rangle\right)=\langle K_{a_{1},\text{f}}(x_{1})K_{a_{2},\text{f}}(x_{2})\mathcal{J}^{(m)}_{\ell,\text{f}}(x_{3},z)\rangle\,,$ (66)
which holds for all four types of bilinear operators. We obtain an equation similar to (2.1):
$\lim_{\epsilon\rightarrow 0}\Bigg{[}$ $\alpha^{-2}4^{2k}\left(\Delta_{\smash{\phi}}-\frac{\Delta_{\mathcal{J}}}{2}+\frac{\ell}{2}\right)_{2k}\left(\Delta_{\smash{\phi}}-\frac{\Delta_{\mathcal{J}}}{2}-\frac{\ell}{2}-\frac{d-2}{2}\right)_{2k}$
×$\times\frac{\lambda_{\phi\phi\mathcal{J}}\left(z\cdot x_{12}\right)^{\ell}\mathbf{t}}{x_{12}^{2\Delta_{\smash{\phi}}+4k-\Delta_{\mathcal{J}}+\ell}|x_{3}|^{2\Delta_{\mathcal{J}}}}+O(\,|x_{3}|^{-2\Delta_{\mathcal{J}}-1})\Bigg{]}=\frac{\lambda_{KK\mathcal{J},\text{f}}\left(z\cdot x_{12}\right)^{\ell}\mathbf{t}}{x_{12}^{2\Delta_{\smash{K_{\text{f}}}}-\Delta_{\mathcal{J}_{\text{f}}}+\ell}|x_{3}|^{2\Delta_{\mathcal{J}_{\text{f}}}}}+O(\,|x_{3}|^{-2\Delta_{\mathcal{J}\smash{,\text{f}}}-1})\,,$ (67)
where $\lambda_{KK\mathcal{J},\text{f}}$ denotes the OPE coefficient of $\mathcal{J}_{ab,\text{f}}\in K_{a,\text{f}}\times K_{b,\text{f}}$. The tensor structures $\mathbf{t}$ on both hand sides match, because the operators on both hand sides of (66) live in the same $S_{N+1}$ representations. The constraint on the anomalous dimensions can be written collectively as
$2^{4k}(2k-m-1)!\,(-m)_{m}(1-3k-\ell-m)_{2k}\lim_{\epsilon\rightarrow 0}\left[\alpha^{-2}\left(\gamma_{\phi}-\frac{\gamma_{\mathcal{J}}}{2}\right)\right]=\frac{\lambda_{KK\mathcal{J},\text{f}}}{\lambda_{\phi\phi\mathcal{J},\text{f}}}\,,$ (68)
which is of the same form as (28). The difference is that here the ratio of free OPE coefficients depend on $N$. The result is
$\gamma_{\mathcal{J}}=2\gamma_{\phi}-\frac{R\,\alpha^{2}}{2^{4k-1}(2k-m-1)!(-m)_{m}(1-3k-\ell-m)_{2k}}+O(\epsilon^{3/2})\,,$ (69)
where the ratio $R\equiv\frac{\lambda_{KK\mathcal{J},\text{f}}}{\lambda_{\phi\phi\mathcal{J},\text{f}}}$ is worked out in appendix E:
$R=\begin{cases}\;4(N-1)(N+1)^{2}\qquad&\text{trivial}\\ \;4(N-2)(N+1)^{2}&\text{standard}\\ \;-4(N+1)^{2}&\text{symmetric traceless}\\ \;-4(N+1)^{2}&\text{antisymmetric}\,,\end{cases}$ (70)
On the highest trajectory $m=k-1$, the spin-2 current in the trivial representation $(N+1)$ is the stress tensor. As expected, the anomalous dimension of the stress tensor is $O(\epsilon^{3/2})$ according to (69). At $k=1$, the set of leading anomalous dimensions $\{\gamma_{\phi},\gamma_{\mathcal{J}}\}$ contains negative elements when $N>7/3$, which indicates that the canonical Potts model is non-unitary near the upper critical dimension. Moreover, there exists a shadow relation in the standard representation:
$\Delta_{\smash{\phi}}+\Delta_{\mathcal{J}_{\text{standard rep}}}\big{|}_{\ell=m=0}=d+O(\epsilon^{3/2})\,,$ (71)
which is true for any positive integer $k$. In the $N\rightarrow\infty$ limit, this equation is related to (30) in the generalized Yang-Lee theory.
Below, we examine the Chew-Frautschi plots to gain further insights into the theory. The analytic continuation of the trajectories is given by Caron-Huot:2022eqs
$(\Delta-d/2)^{2}$
=\;= $\left(d-2k+2m+\ell+\gamma_{\mathcal{J}}-d/2\right)^{2}$
=\;= $\left(k+2m+\ell+2\gamma_{1}-\frac{R\,\alpha^{2}}{2^{4k-1}(2k-m-1)!(1-3k-m-\ell)_{2k}(-m)_{m}}-\frac{\epsilon}{2}+O(\epsilon^{3/2})\right)^{2}\,.$ (72)
Using this relation between $\Delta$ and $\ell$, we plot the Regge trajectories of the bilinear operators $\mathcal{J}^{(m)}_{\ell}$.
Figure 1: Chew-Frautschi plot for $\mathcal{J}^{(m=0)}_{\ell,\,\text{trivial rep}}$ at $N=3$. The dashed lines indicate the the $\epsilon\rightarrow 0$ limit of the Regge trajectory. We have set $k=1$ and $\epsilon=10^{-3}$.
Figure 2: Regge trajectories of $\mathcal{J}^{(m=0)}_{\ell\,\text{trivial rep}}$. The $N=\infty$ case corresponds to the decoupled Yang-Lee theory. The plot is made at $k=1$ and $\epsilon=10^{-3}$.
Figure 3: Regge trajectories associated with $\mathcal{J}^{(m=0)}_{\ell,\,\text{trivial rep}}$. The trajectory in the interacting theory approaches the free theory trajectory as $N\rightarrow 1$. We have used $k=1$ and $\epsilon=10^{-3}$.
We first focus on the canonical case $k=1$. A Chew-Frautschi plot associated with bilinear operators in the trivial representation is given in figure 1. The plot varies with $N$, as shown in figure 2. The trajectories intersect with the horizontal axis. Substituting $\ell=0$ into (3.1), we find two solutions
$\Delta^{*}_{\pm}-d/2=\pm\left(1+\frac{3R^{\prime}-11N+23}{6(3N-7)}\,\epsilon+O(\epsilon^{3/2})\right)\,,$ (73)
where we have defined $R^{\prime}\equiv R/(N+1)^{2}$. The solution $\Delta^{*}_{-}$ in the standard representation is precisely the scaling dimension of $\phi_{a}$, and $\Delta^{*}_{+}$ is given by $\Delta_{\mathcal{J_{\text{standard rep}}}}|_{\ell=m=0}$. This is consistent with the shadow relation (71), since $\Delta^{*}_{\pm}$ are related by the shadow transformation $\Delta\rightarrow d-\Delta$.
Figure 4: An $\mathbb{R}^{3}$ projection of the $\mathbb{C}^{2}$ Chew-Frautschi plot for $\mathcal{J}^{(m=0)}_{\ell,\,\text{trivial rep}}$ at $N=3$. The surfaces are the Riemann surfaces of the functions $\ell(\Delta)$. The imaginary part of $\ell$ is shown by color. Red regions correspond to $\text{Im}[\ell]>0$, and blue regions are associated with $\text{Im}[\ell]<0$. The two turquoise points represent two solutions to $\ell=0$, which are $\Delta_{\pm}$ in (73). The plot is symmetric with respect to the plane $\text{Re}[\Delta-d/2]=0$. We have chosen the parameters $k=1$ and $\epsilon=0.3$.
Figure 5: An $\mathbb{R}^{3}$ projection of the $\mathbb{C}^{2}$ Regge trajectories of $\mathcal{J}^{(m=0)}_{\ell,\,\text{trivial rep}}$ at $N=3$. The peaks at $\text{Re}[\ell]=-2$ and $\text{Im}[\ell]=0$ correspond to the $1/(\ell+2)$ pole in the anomalous dimension. We indicate the imaginary part of $\Delta-d/2$ by color. The regions with $\text{Im}[\Delta-d/2]>0$ are in red, and the regions with $\text{Im}[\Delta-d/2]<0$ are in blue. Two of the solutions to $\Delta=d/2$ are $\ell_{\pm}$ in (74). They form a complex conjugate pair, and are indicated by the two turquoise points. We have used $k=1$ and $\epsilon=0.3$.
Figure 6: A 3D Chew-Frautschi plot for $\mathcal{J}^{(m=0)}_{\ell,\,\text{trivial rep}}$ at $N=6/5$. Color coding indicates the imaginary part of $\ell$: red for $\text{Im}[\ell]>0$ and blue for $\text{Im}[\ell]<0$. The turquoise points are the solutions to $\ell=0$, given by $\Delta_{\pm}$ in (73). The plot is symmetric relative to the plane $\text{Re}[\Delta-d/2]=0$. The parameters are $k=1$ and $\epsilon=0.3$.
(a) A view showing $\ell_{-}^{*}$.
(b) A view showing $\ell_{+}^{*}$.
Figure 7: An $\mathbb{R}^{3}$ projection of the $\mathbb{C}^{2}$ Chew-Frautschi plot for $\mathcal{J}^{(m=0)}_{\ell,\,\text{trivial rep}}$ at $N=6/5$. The peaks at $\text{Re}[\ell]=-2$, $\text{Im}[\ell]=0$ is due to the $1/(\ell+2)$ pole in the anomalous dimension. The imaginary part of $\ell$ is shown by color: red regions implies $\text{Im}[\ell]>0$, and blue regions indicate $\text{Im}[\ell]<0$. The turquoise points mark the solutions $\ell_{\pm}^{*}$ in (74). The plot is made at $k=1$ and $\epsilon=0.3$.
There are also intersections with the vertical axis. We substitute $\Delta=d/2$ into (3.1) and obtain three solutions
$\ell^{*}_{\pm}$ $=-1\pm i\,\sqrt{\frac{2R^{\prime}}{3N-7}}\;\epsilon^{1/2}+O(\epsilon)\,,$ (74)
$\ell^{*}_{3}$ $=-2-\frac{2R^{\prime}}{3N-7}\,\epsilon+\frac{2R^{\prime}(12R^{\prime}+11N-23)}{3(3N-7)^{2}}\epsilon^{2}+O(\epsilon^{3})\,.$ (75)
Let us discuss the case of the trivial representation, where $R^{\prime}=4(N-1)$. For $N>7/3$, the solutions $\ell^{*}_{\pm}$ correspond to complex spins, so they do not show up in the real $(\Delta-d/2,\ell)$ plane. For $31<N<7/$ 3, there are three intersection points, as in figure 3, since $R^{\prime}$ is negative and $\ell^{*}_{\pm}$ are real. At $N=1$, we have $R^{\prime}=0$ and $\gamma_{1}=O(\epsilon^{3/2})$, meaning that the anomalous dimensions are zero to order $\epsilon^{1}$. So the $N=1$ trajectory reduces to that in the free theory. [^20] Strictly speaking, the free trajectory should not include the horizontal line $\ell=-2$. The $N<1$ plots are similar to the $N>7/3$ ones, where $\ell^{*}_{\pm}$ are complex. Furthermore, we show the projections of $\mathbb{C}^{2}$ Chew-Frautschi plots in figures 4, 5, 6 and 7. For the standard representation, $R^{\prime}$ changes sign at $N=2$. There is one intersection point with the vertical axis for $N>7/3$ as well as for $N<2$, and there are three intersection points for $32<N<7/$ 3. For the remaining two representations, $R^{\prime}$ is always negative. There are only two scenarios: $N>7/3$ and $N<7/3$. The former gives the trajectories with three intersection points on the vertical axis, and the latter yields the trajectories with one intersection point.
Now, we briefly comment on the $k>1$ Chew-Frautschi plots. One difference is the presence of more curves, which arise from the deformation of more horizontal lines in the Gaussian limit. Another difference is that we can consider trajectories with higher $m$. Let us first discuss the Gaussian limit, where the scaling dimensions do not depend on the $S_{N+1}$ representations. For a positive integer $k$, the horizontal lines are given by $\frac{(1-3k-\ell-m)_{2k}}{\ell+k+2m}=0$, implying that there are $12k-$ 1 such lines. The diagonal lines are translations of those in the $k=1$ case. In the interacting theory, the trajectories are deformed into disconnected pieces. For example, the $\mathcal{J}^{(m=0)}_{\ell,\,\text{trivial rep}}$ trajectory in the $k=2$ theory and its Gaussian limit are shown in figure 8. As in the $k=1$ case, the Chew-Frautschi plots in the interacting theory depend on $N$. For both $m=0$ and $m=1$, there are three scenarios: $N>65/37$, $371<N<65/$ 37 and $N<1$. We present the Chew-Frautschi plots with different values of $N$ in figures 9, 10, 11 and 12. The Chew-Frautschi plots with $N<1$ are similar to case of $N>65/37$.
Figure 8: Chew-Frautschi plot for $\mathcal{J}^{(m=0)}_{\ell,\,\text{trivial rep}}$ in the $k=2$ theory. The interacting case with $N=2$ is shown in red. In the Gaussian limit, there are two dashed diagonal lines given by $(\ell+2)^{2}=(\Delta-6)^{2}$, and there exist three dashed horizontal lines at $\ell=-3,-4,-5$.
Figure 9: In the $k=2$ theory, the Regge trajectories of $\mathcal{J}^{(m=0)}_{\ell}$ are shown. These trajectories correspond to the case of $N>65/37$. We have set $\epsilon=10^{-3}$.
Figure 10: Chew-Frautschi plot of the $k=2$ interacting theory. The Regge trajectories of $\mathcal{J}^{(m=0)}_{\ell,\,\text{trivial rep}}$ with $371<N<65/$ 37 are shown. We have used $\epsilon=10^{-3}$.
Figure 11: Regge trajectories of $\mathcal{J}^{(m=1)}_{\ell,\,\text{trivial rep}}$ in the $k=2$ interacting theory. We present the scenario where $N>65/37$. Again, $\epsilon=10^{-3}$ is used.
Figure 12: In the $k=2$ interacting theory, we show the Regge trajectories of $\mathcal{J}^{(m=1)}_{\ell,\,\text{trivial rep}}$ with $371<N<65/$ 37, where $\epsilon$ is set to $10^{-3}$.
3.2 Analytic bootstrap
In this subsection, we use the analytic bootstrap to check the results from the multiplet recombination. As in subsection 2.2, we consider the crossing constraint on the four-point function
$\langle\phi_{a}(x_{1})\phi_{b}(x_{2})\phi_{c}(x_{3})\phi_{d}(x_{4})\rangle=\frac{g_{abcd}(z,\bar{z})}{x_{12}^{2\Delta_{\smash{\phi}}}x_{34}^{2\Delta_{\smash{\phi}}}}\,.$ (76)
The tensor structures of the four-point function are given by (138), (139), (140) and (141) in appendix D. We consider
$g_{abcd}(z,\bar{z})=\sum_{i=1}^{4}g^{(i)}(z,\bar{z})\,t_{abcd}^{(i)}\,.$ (77)
Then, the crossing equations are written as (D), (D), (D) and (D). The functions $g^{(i)}(z,\bar{z})$ decompose into conformal blocks associated with the operators in the trivial, standard, symmetric traceless and antisymmetric representations. As in subsection 2.2, we focus on the leading order in the $z\rightarrow 0$ expansion, and we verify the multiplet-recombination results for the $m=0$ trajectory.
At order $\epsilon^{0}$, the identity operator contributes to the $\bar{z}\rightarrow 1$ enhanced singularity on the right-hand sides of the crossing equations. Since the identity operator lives in the trivial representation, the four types of free squared OPE coefficients are
$P_{\mathcal{J}^{(0)}_{\ell},\text{f}}^{\text{ generic }N}=P_{\mathcal{J}^{(0)}_{\ell},\text{f}}^{\;N=1}\times\begin{cases}\;\frac{1}{N}\qquad&\text{trivial}\\ \;\frac{1}{(N-1)(N+1)^{2}}&\text{standard}\\ \;1&\text{symmetric traceless}\\ \;-1&\text{antisymmetric}\,,\end{cases}$ (78)
[…]
At order $\epsilon^{1}$, the enhanced singularities on the right-hand sides of the crossing equations are given by the operator $\phi_{a}$ in the standard representation. Then, the anomalous dimensions are related to those in the Yang-Lee case by
$\tilde{\gamma}_{\mathcal{J}}\big{|}_{m=0}=\frac{P_{\phi}^{\text{Potts}}}{P_{\phi}^{\text{YL}}}\tilde{\gamma}_{\mathcal{J}}^{\text{YL}}\big{|}_{m=0}\times\begin{cases}\;\frac{(N-1)(N+1)^{2}}{N}\times N\qquad&\text{trivial}\\ \;\frac{N-2}{N-1}\times(N-1)(N+1)^{2}&\text{standard}\\ \;-(N+1)^{2}\times 1&\text{symmetric traceless}\\ \;(N+1)^{2}\times(-1)&\text{antisymmetric}\,,\end{cases}$ (79)
where $P_{\phi}^{\text{YL}}$ and $\tilde{\gamma}_{\mathcal{J}}^{\text{YL}}$ are the Yang-Lee data. This means that we consider the matching of the lightcone singularity in the Yang-Lee case, then we add the $N$-dependent factors and replace the Yang-Lee OPE coefficients with those in the Potts model. Using the OPE coefficients (24) and (63), [^22] According to (37), the OPE coefficients with different normalizations are related by $P_{\phi}^{\text{YL}}=\left(\lambda_{\phi\phi\phi}^{\text{YL}}\right)^{2}$ and $P_{\phi}=\lambda_{\phi\phi\phi}^{2}$. we verify that the result agrees with (69) from the multiplet recombination. We have also examined the $m=1$ trajectory in appendix B, and we confirm that the multiplet-recombination results are consistent with crossing symmetry.
4 Generalized $\phi^{2n+1}$ theory ($n>1$)
In this section, we study the $\phi^{2n+1}$ deformation of the free higher-derivative theory, with the integer $n>1$. To avoid derivative interactions, we require that $k$ and $12n-$ 1 have no common divisor. [^23] This is similar to the case of the $\phi^{2n}$ theory, where $k$ and $n-1$ have no common divisor Safari:2017irw ; Safari:2017tgs . The conditions for the $\phi^{2n+1}$ and $\phi^{2n}$ theories can be unified as follows. For a $\phi^{p}$ theory with integer $p$, we require that $k$ and $p-2$ have no common divisor other than 2. In this way, we can consistently consider the pure $\phi^{2n+1}$ interaction. The action at $d=d_{\text{u}}-\epsilon$ takes the form
$S=\int\mathrm{d}^{d}x\left(\phi\,\Box^{k}\phi+g\mu^{(2n-1)\epsilon/2}\phi^{2n+1}\right)\,,$ (80)
where the upper critical dimension is given by
$d_{\text{u}}=2k\,\frac{2n+1}{2n-1}\,.$ (81)
This theory describes the generalized multicritical Yang-Lee edge singularity vonGehlen:1994rp ; Lencses:2022ira ; Lencses:2023evr ; Lencses:2024wib . For $k=1$, it was argued that the Lagrangian in (80) can be obtained from the Landau-Ginzburg Lagrangian of the multicritical Ising model. For the latter Lagrangian, a shift of the order parameter field by a purely imaginary constant generates $\mathbb{Z}_{2}$-odd terms. Then, one can tune the coupling constants to reach the multicritical point given by (80). [^24] Not all multicritical Ising models lead to the Lagrangian in (80). See Lencses:2024wib for more details. This generalizes Fisher’s argument in Fisher:1978pf .
In this section, the CFT data is not completely fixed by the multiplet recombination. We determine the set of anomalous dimensions up to one unknown parameter $\alpha$, which can be fixed using the diagrammatic method in the canonical case. For arbitrary $\alpha$, We also check the consistency of the results with crossing symmetry using the analytic bootstrap.
4.1 Multiplet recombination
The generalization of (6) reads
$\lim_{\epsilon\rightarrow 0}\left(\alpha^{-1}\Box^{k}\phi\right)=\phi^{2n}_{\text{f}}\,.$ (82)
For $n>1$, the normalization $\alpha$ is not determined by the matching conditions generated from this equation, and $\alpha$ appears in the results for the anomalous dimensions, as well as the OPE coefficients. This distinction from the $n=1$ case is due to the vanishing of free correlator $\langle\phi^{2n}_{\text{f}}(x_{1})\phi_{\text{f}}(x_{2})\phi_{\text{f}}(x_{3})\rangle$ with $n>1$. In subsection 2.1, we see that the free theory data is an important input in the multiplet recombination approach. As there is less free theory input available here, the set of matching conditions is less constraining. Nevertheless, we can still obtain a consistent set of data using (82).
For two-point functions, the matching condition takes the form
$\lim_{\epsilon\rightarrow 0}\left(\alpha^{-2}\langle\Box^{k}\phi(x_{1})\Box^{k}\phi(x_{2})\rangle\right)=\langle\phi^{2n}_{\text{f}}(x_{1})\phi^{2n}_{\text{f}}(x_{2})\rangle\,.$ (83)
The left-hand side has the same form as that of (7). The right-hand side is a free correlator which can be computed using Wick’s theorem. The result is
$\lim_{\epsilon\rightarrow 0}\left[\alpha^{-2}\,4^{2k}\left(\Delta_{\smash{\phi}}\right)_{2k}\left(\Delta_{\smash{\phi}}-\frac{d-2}{2}\right)_{2k}\frac{1}{x_{12}^{2\Delta_{\smash{\phi}}+4k}}\,\right]=\frac{(2n)!}{x_{12}^{2\Delta_{\smash{\phi^{2n}_{\text{f}}}}}}\,.$ (84)
The functional forms match, since $\lim_{\epsilon\rightarrow 0}(2\Delta_{\smash{\phi}}+4k)=2\Delta_{\smash{\phi^{2n}_{\text{f}}}}$. The definition of the anomalous dimension is
$\gamma_{\phi}=\Delta_{\smash{\phi}}-\frac{d-2k}{2}\,,\qquad d=d_{\text{u}}-\epsilon\,.$ (85)
The constraint on the anomalous dimension reads
$4^{2k}k!\left(\frac{d_{\text{u}}-2k}{2}\right)_{2k}(1-k)_{k-1}\lim_{\epsilon\rightarrow 0}\left(\alpha^{-2}\gamma_{\phi}\right)=(2n)!\,,$ (86)
where the upper critical dimension $d_{\text{u}}$ is given by (81). We obtain a relation between the anomalous dimension $\gamma_{\phi}$ and the unknown parameter $\alpha$:
$\gamma_{\phi}=\frac{(2n)!\,\alpha^{2}}{4^{2k}k!\left(\frac{2k}{2n-1}\right)_{2k}(1-k)_{k-1}}+O(\alpha^{3})\,.$ (87)
The next step is to consider three-point functions. Unfortunately, the matching condition for $\langle\Box^{k}\phi(x_{1})\phi(x_{2})\phi(x_{3})\rangle$ is not useful, as the corresponding free correlator $\langle\phi^{2n}_{\text{f}}(x_{1})\phi_{\text{f}}(x_{2})$ $\phi_{\text{f}}(x_{3})\rangle$ vanishes for $n>1$. Moreover, the matching condition for $\langle\Box^{k}\phi(x_{1})\Box^{k}\phi(x_{2})\Box^{k}\phi(x_{3})\rangle$ involves $\alpha$ and an unknown OPE coefficient $\lambda_{\phi\phi\phi}$. So we cannot determine $\alpha$ and $\gamma_{\phi}$ by following the procedure in section 2.1. Nevertheless, we cam still find the relation between $\lambda_{\phi\phi\phi}$ and $\alpha$ using the matching condition
$\lim_{\epsilon\rightarrow 0}\left(\alpha^{-3}\langle\Box^{k}\phi(x_{1})\Box^{k}\phi(x_{2})\Box^{k}\phi(x_{3})\rangle\right)=\langle\phi^{2n}_{\text{f}}(x_{1})\phi^{2n}_{\text{f}}(x_{2})\phi^{2n}_{\text{f}}(x_{3})\rangle\,.$ (88)
The left-hand side is computed in appendix A. The result takes the same form as the left-hand side of (2.1). We obtain the constraint on the OPE coefficient:
$(-1)^{k}4^{3k}\left(\frac{d_{\text{u}}-6k}{4}\right)_{3k}\left(\frac{d_{\text{u}}-2k}{4}\right)_{k}^{3}\;\lim_{\epsilon\rightarrow 0}\left(\alpha^{-3}\lambda_{\phi\phi\phi}\right)=\left(\frac{(2n)!}{n!}\right)^{3}\,.$ (89)
The solution is
$\lambda_{\phi\phi\phi}=\frac{(-1)^{k}\left((2n)!\right)^{3}\alpha^{3}}{4^{3k}\left(n!\right)^{3}\left(\frac{k}{2n-1}\right)_{k}^{3}\left(-2k\frac{n-1}{2n-1}\right)_{3k}}+O(\alpha^{3})\,,$ (90)
which agrees with Codello:2017qek at $k=1$.
To obtain more useful constraints on the anomalous dimensions, we should avoid the matching conditions where the left-hand sides involve vanishing OPE coefficients in the Gaussian limit. In this way, we avoid introducing more unknown OPE coefficients as we examine more correlation functions. Additionally, the free correlators on the right-hand sides of the matching conditions should not be zero. Since (82) changes the $\mathbb{Z}_{2}$ parity under $\phi\rightarrow-\phi$, we have to consider the matching conditions with two $\Box^{k}\phi$’s. An example is
$\lim_{\epsilon\rightarrow 0}\left(\alpha^{-2}\langle\Box^{k}\phi(x_{1})\Box^{k}\phi(x_{2})\phi^{2}(x_{3})\rangle\right)=\langle\phi^{2n}_{\text{f}}(x_{1})\phi^{2n}_{\text{f}}(x_{2})\phi^{2}(x_{3})\rangle\,,$ (91)
which yields a relation between the anomalous dimensions of $\phi$ and $\phi^{2}$. This is a special case of the matching conditions for bilinear operators.
Let us consider the more general case. The matching condition involving a primary bilinear operator $\mathcal{J}^{(m)}_{\ell}$ reads
$\lim_{\epsilon\rightarrow 0}\left(\alpha^{-2}\langle\Box^{k}\phi(x_{1})\Box^{k}\phi(x_{2})\mathcal{J}^{(m)}_{\ell}(x_{3},z)\rangle\right)=\langle\phi^{2n}_{\text{f}}(x_{1})\phi^{2n}_{\text{f}}(x_{2})\mathcal{J}^{(m)}_{\ell}(x_{3},z)\rangle\,.$ (92)
As in (2.1), the constraint is extracted from the leading term in the $x_{3}\rightarrow\infty$ expansion. We have
$\lim_{\epsilon\rightarrow 0}\Bigg{[}$ $\alpha^{-2}4^{2k}\left(\Delta_{\smash{\phi}}-\frac{\Delta_{\mathcal{J}}}{2}+\frac{\ell}{2}\right)_{2k}\left(\Delta_{\smash{\phi}}-\frac{\Delta_{\mathcal{J}}}{2}-\frac{\ell}{2}-\frac{d-2}{2}\right)_{2k}$
×$\times\frac{\lambda_{\phi\phi\mathcal{J}}\left(z\cdot x_{12}\right)^{\ell}}{x_{12}^{2\Delta_{\smash{\phi}}+4k-\Delta_{\mathcal{J}}+\ell}|x_{3}|^{2\Delta_{\mathcal{J}}}}+O(\,|x_{3}|^{-2\Delta_{\mathcal{J}}-1})\Bigg{]}=\frac{\lambda_{\phi^{2n}\phi^{2n}\mathcal{J},\text{f}}\left(z\cdot x_{12}\right)^{\ell}}{x_{12}^{2\Delta_{\smash{\phi^{2n}_{\text{f}}}}-\Delta_{\mathcal{J}\smash{,\text{f}}}+\ell}|x_{3}|^{2\Delta_{\mathcal{J},\text{f}}}}+O(\,|x_{3}|^{-2\Delta_{\mathcal{J}\smash{,\text{f}}}-1})\,.$ (93)
Again, one can check that the functional forms match using $\lim_{\epsilon\rightarrow 0}(2\Delta_{\smash{\phi}}+4k)=2\Delta_{\smash{\phi^{2n}_{\text{f}}}}$. According to the definition of the anomalous dimension $\gamma_{\mathcal{J}}=\Delta_{\mathcal{J}}-(d-2k+2m+\ell)$, the equation above leads to
$4^{2k}(2k-m-1)!\,(-m)_{m}\left(1-\frac{d_{\text{u}}}{2}-\ell-m\right)_{2k}\lim_{\epsilon\rightarrow 0}\left[\alpha^{-2}\left(\gamma_{\phi}-\frac{\gamma_{\mathcal{J}}}{2}\right)\right]=\frac{\lambda_{\phi^{2n}\phi^{2n}\mathcal{J},\text{f}}}{\lambda_{\phi\phi\mathcal{J},\text{f}}}\,,$ (94)
where the ratio of OPE coefficients $\frac{\lambda_{\phi^{2n}\phi^{2n}\mathcal{J},\text{f}}}{\lambda_{\phi\phi\mathcal{J},\text{f}}}=2n(2n)!$ is computed in appendix E. The solution for the anomalous dimension of the bilinear operator is
$\gamma_{\mathcal{J}}=2\gamma_{\phi}-\frac{n(2n)!\,\alpha^{2}}{2^{4k-2}(2k-m-1)!(-m)_{m}\left(1-k\frac{2n+1}{2n-1}-\ell-m\right)_{2k}}+O(\alpha^{3})\,,$ (95)
where $\gamma_{\phi}$ is given in (87). The special case (91) is associated with $\ell=m=0$. One can check that, to order $\alpha^{2}$, the anomalous dimension of the stress tensor $(m=k-1,\;\ell=2)$ vanishes, even though we do not know the explicit expression for $\alpha$. In other words, we cannot determine $\alpha$ by requiring that the stress tensor does not acquire anomalous dimension. [^25] We believe that the spinning recombination cannot fix $\alpha$ either. For example, at $k=1$, we find that $\lim_{\epsilon\rightarrow 0}(\alpha^{-2}\langle\Box\phi(x_{1})\phi(x_{2})\partial\mathcal{J}(x_{3}))\rangle$ is proportional to $\lim_{\epsilon\rightarrow 0}(\alpha^{-2}\gamma_{\mathcal{J}})$, so the matching condition should only yield a relation between $\gamma_{\mathcal{J}}$ and $\alpha$, and this relation ought to be the same as (95).
Since the anomalous dimensions are not fixed by our CFT analysis, we resort to inputs from the diagrammatic method. At $k=1$, Codello et al. computed the anomalous dimensions of some scalar operators in the $n>1$ theory Codello:2017epp . The result for $\gamma_{\phi}$ is expressed in terms of the critical coupling constant in their equation (5.7), which is consistent with our formula (87). Furthermore, they also obtained the critical coupling constant, so they fully determined the anomalous dimension $\gamma_{\phi}$. We fix $\alpha$ using their results and find [^26] In Codello:2017epp , the general expression for the critical coupling constant $g$ in equation (5.6) may contain a typo, as it does not seem to reproduce their $n=2$ result in equation (3.2). Here we use their beta functionals to solve for $g$, and the result is consistent with their equation (3.2).
$\alpha^{2}=\frac{96\epsilon}{(2n-1)(2n)!\left(6-(2n+1)!(2n)!\,\Gamma\left(\frac{2n+1}{2n-1}\right)\mathcal{S}(n)\right)}+O(\epsilon^{2})\,,$ (96)
which involves a sum
$\mathcal{S}(n)=\sum_{\begin{subarray}{c}r+s+t=2n+1\\ r,s,t\in\mathbb{Z}_{+}\end{subarray}}\Bigg{[}$ $\;\frac{1}{r!s!t!(2n+1-r-s)!(2n+1-s-t)!(2n+1-t-r)!}$
×$\quad\times\frac{\Gamma\left(\frac{s+t-r}{2n-1}\right)\Gamma\left(\frac{t+r-s}{2n-1}\right)\Gamma\left(\frac{r+s-t}{2n-1}\right)}{\Gamma\left(\frac{2r}{2n-1}\right)\Gamma\left(\frac{2s}{2n-1}\right)\Gamma\left(\frac{2t}{2n-1}\right)}\;\Bigg{]}\,.$ (97)
Here $\mathbb{Z}_{+}$ denotes the set of positive integers. If we use this input from the diagrammatic method, then the anomalous dimensions of $\phi$ and $\mathcal{J}$ are completely fixed for $k=1$. The $n=1$ case of (96) matches the Yang-Lee result (23). For $n=2,3,4$, our results for $\gamma_{\phi}$ are in agreement with those of Gracey:2017okb . [^27] However, our results for $\gamma_{\phi^{2}}$ do not agree with Gracey:2017okb . Some disagreement was pointed out by Codello:2017epp as well, and it was shown that the results of Codello:2017epp are supported by the CFT analysis Codello:2017qek .
4.2 Analytic bootstrap
In this subsection, we verify the multiplet-recombination results using the analytic bootstrap. For arbitrary $\alpha$, the multiplet-recombination results are consistent with crossing symmetry, according to the analytic bootstrap in this section. As in subsection 2.2, we consider the four-point function
$\langle\phi(x_{1})\phi(x_{2})\phi(x_{3})\phi(x_{4})\rangle=\frac{g(z,\bar{z})}{x_{12}^{2\Delta_{\smash{\phi}}}x_{34}^{2\Delta_{\smash{\phi}}}}\,,$ (98)
and the crossing equation
$g(z,\bar{z})=\frac{z^{\Delta_{\smash{\phi}}}\bar{z}^{\Delta_{\smash{\phi}}}}{(1-z)^{\Delta_{\smash{\phi}}}(1-\bar{z})^{\Delta_{\smash{\phi}}}}g(1-\bar{z},1-z)\,.$ (99)
At order $\alpha^{0}$, this analysis is the similar to that in subsection 2.2. The difference is that the scaling dimension of $\phi_{\text{f}}$ takes a different value $\Delta_{\smash{\phi_{\text{f}}}}=\frac{2k}{2n-1}$. We derive the free OPE coefficient
$P_{\smash{\mathcal{J}_{\ell}^{(0)}},\text{f}}=2S_{\frac{2k}{1-2n}}(h)=2\frac{\left(\frac{2k}{2n-1}\right)_{\ell}^{2}}{\ell!\left(\frac{4k}{2n-1}+\ell-1\right)_{\ell}}\,,$ (100)
where we have used the relation (39) and the formula (40).
$\phi$$\phi$$\phi$$\phi$$\phi^{2n-1}$$g$$g$
Figure 13: Feynman diagram in the Lagrangian description (80). Due to the $g\phi^{2n+1}$ vertices, the operator $\phi^{2n-1}$ contributes at order $g^{2}\sim\alpha^{2}$.
At order $\alpha^{2}$, the scalar $\phi^{2n-1}$ contributes to the enhanced singularity on the right-hand side of the crossing equation, due to the squared OPE coefficient $P_{\phi^{2n-1}}\sim\alpha^{2}$ (see figure 13). The $\log z$ part of the $\phi^{2n-1}$ block is given by (44), with $\mathcal{O}=\phi^{2n-1}$. The corresponding term takes the form
$\frac{\bar{z}^{\Delta_{\smash{\phi}}}}{(1-\bar{z})^{\Delta_{\smash{\phi}}}}P_{\phi^{2n-1}}G_{\phi^{2n-1}}(1-\bar{z},1-z)\Big{|}_{\log z\text{ part}}$
=\;= $\frac{n(2n)!\,\Gamma(2k)\alpha^{2}}{2^{4k-1}\Gamma(k)^{2}(1-2k)_{k}^{2}\left(k\frac{3-2n}{2n-1}\right)_{k}^{2}}\,\bar{z}^{\frac{2k}{2n-1}}(1-\bar{z})^{k\frac{2n-3}{2n-1}}{}_{2}F_{1}\!\left(k,k;k\frac{2n-3}{2n-1}+1;1-\bar{z}\right)+O(\alpha^{3})\,,$ (101)
where we have used the squared OPE coefficient (160) derived in appendix F. We should be more careful, as there are infinitely many terms contributing to the enhanced singularity. A more elegant approach is to use the Lorentzian inversion formula. In the lightcone limit, this leads to an SL($2,\mathbb{R}$) inversion integral Caron-Huot:2017vep ; Alday:2017zzv
$C(z,h)=\frac{(2h-1)\Gamma(h)^{4}}{\pi^{2}\Gamma(2h)^{2}}\int_{0}^{1}\frac{\mathrm{d}\bar{z}}{\bar{z}^{2}}k_{h}(\bar{z})\,\text{dDisc}[g(1-\bar{z},1-z)]\Big{|}_{\text{leading term in $z$}}\,,$ (102)
where we consider the leading term in the $z\rightarrow 0$ expansion. The double discontinuity is defined as
$\text{dDisc}[f(\bar{z})]=f(\bar{z})-\frac{1}{2}f^{\circlearrowleft}(\bar{z})-\frac{1}{2}f^{\circlearrowright}(\bar{z})\,,$ (103)
where $f^{\circlearrowleft}$ and $f^{\circlearrowright}$ are two analytic continuations around $\bar{z}=1$. The result of the integral is expanded as
$C(z,h)=z^{\Delta_{\phi}}P_{\mathcal{J}}\,\left(1+\frac{1}{2}\tilde{\gamma}_{\mathcal{J}}\log z+O(\tilde{\gamma}^{2}_{\mathcal{J}})\right)+O(z^{\Delta_{\phi}+1})\,.$ (104)
The calculation of the dDisc associated with (4.2) amounts to multiplying by the factor $2\sin^{2}(\pi k\,\frac{2n-3}{2n-1})$. Substituting this dDisc into the inversion integral (102) and dividing by the free OPE coefficient (100), we obtain the anomalous dimensions of the bilinear operators on the $m=0$ trajectory. The results agree with those from the multiplet recombination, confirming that the multiplet-recombination results are crossing symmetric. In appendix B, the analytic bootstrap has also been carried out at subleading twist, where the multiplet-recombination results are again consistent with crossing symmetry.
5 Discussion
In this paper we have computed the anomalous dimensions of $\phi$ and bilinear operators $\mathcal{J}$ in the generalized $\phi^{2n+1}$ theory, using the conformal multiplet recombination. Our main results are summarized in table 1. We have verified that these results are consistent with crossing symmetry using the analytic bootstrap.
We have considered the $S_{N+1}$ global symmetry, but have restricted to the $\phi^{3}$ case. It might be interesting to study the $S_{N+1}$-symmetric $\phi^{p}$ theories with higher $p$, as they should provide a variety of universality classes. In particular, the canonical $\phi^{4}$ and $\phi^{5}$ theories with $S_{N+1}$ symmetry have been studied in Rong:2017cow ; Codello:2018nbe ; Codello:2020mnt . These theories correspond to the quartic and quintic Potts models. It would also be interesting to investigate their higher-derivative generalizations.
Since the set of anomalous dimensions is fixed up to an unknown parameter $\alpha$ in the $n>1$ theory, another important question is if we can determine $\alpha$ in the $n>1$ theory by conformal symmetry and consistency requirements. In the canonical case $k=1$, we fix the value of $\alpha$ by the input from Codello:2017epp , which is given in (96). Given the complicated summation in (4.1), it seems unlikely that we can determine $\alpha$ by the leading analysis of the multiplet recombination method. As (4.1) shares some similarities with the subleading corrections in the $\phi^{2n}$ theory Codello:2017epp , we expect that $\alpha$ can be determined by a higher order study, which requires a merging of the multiplet recombination and crossing analyses. As a necessary step, we verified that the multiplet-recombination results are consistent with crossing symmetry for the leading corrections.
As the generalized free theory is non-unitary, we expect that the $\phi^{2n+1}$ deformations are non-unitary as well. The nonperturbative bootstrap study of their integer $d$ versions requires bootstrap methods that do not rely on positivity constraints Gliozzi:2013ysa ; Gliozzi:2014jsa ; El-Showk:2016mxr ; Esterlis:2016psv ; Li:2017agi ; Li:2017ukc ; Li:2021uki ; Kantor:2021kbx ; Kantor:2021jpz ; Afkhami-Jeddi:2021iuw ; Laio:2022ayq ; Li:2023tic . In Guo:2023qtt and the present work, we have investigated the $\phi^{2n}$ and $\phi^{2n+1}$ deformations of the generalized free theory, i.e., $\phi^{m}$ interactions with integer power $m\geq 3$. A natural extension is to consider non-integer $m$. Some simple examples are the fractional power $m=\frac{m_{1}}{2},\frac{m_{1}}{3},\frac{m_{1}}{4},\dots$, where $m_{1}$ is a large enough integer. The $\phi^{m}$ interaction with irrational power or even complex power are also interesting from the perspective of analytic continuation, as they can connect the standard integer-power cases in the theory space, but we expect that they are non-unitary even for the canonical case of $k=1$. The $\mathcal{PT}$ symmetric theories are particularly relevant because they can satisfy the physical requirement of a bounded-from-below spectrum in a continuous range of power Bender:1998ke ; Bender:2007nj ; Bender:2023cem . Recently, the low dimensional $\mathcal{PT}$ symmetric models have been revisited by bootstrap methods in Li:2022prn ; Khan:2022uyz ; Li:2023nip ; John:2023him ; Li:2023ewe ; Li:2024rod , using the equations of motion in the Lagrangian or Hamiltonian formalism. It would be interesting to extend the low dimensional bootstrap methods to the higher dimensional cases. Along the lines of analytic continuation, one may also bootstrap the non-perturbative conformal field theory by the homotopy continuation method Li:2024ggr , where the $\epsilon$ expansion can provide regular boundary conditions near but away from the singular Gaussian limit.
Acknowledgments
This work was supported by the Natural Science Foundation of China (Grant No. 12205386) and the Guangzhou Municipal Science and Technology Project (Grant No. 2023A04J0006).
Appendix A The action of Laplacians
In this appendix, we compute the action of Laplacians on a generic scalar three-point function
$\langle\mathcal{O}_{1}(x_{1})\mathcal{O}_{2}(x_{2})\mathcal{O}_{3}(x_{3})\rangle\propto\frac{1}{x_{12}^{h_{123}}x_{13}^{h_{132}}x_{23}^{h_{231}}}\,,$ (105)
where $h_{ijk}$ are defined by
$h_{ijk}=\Delta_{i}+\Delta_{j}-\Delta_{k}\,.$ (106)
Here $\Delta_{i}$ denotes scaling dimension of $\mathcal{O}_{i}$. The result of the $\Box_{x_{1}}$ action takes the form
$\Box_{x_{1}}\frac{1}{x_{12}^{h_{123}}x_{13}^{h_{132}}x_{23}^{h_{231}}}$
=\;= $\left[\frac{h_{123}\left(h_{123}+h_{132}+2-d\right)}{x_{12}^{2}}+\frac{h_{132}\left(h_{123}+h_{132}+2-d\right)}{x_{13}^{2}}-\frac{h_{123}h_{132}\,x_{23}^{2}}{x_{12}^{2}x_{13}^{2}}\right]\frac{1}{x_{12}^{h_{123}}x_{13}^{h_{132}}x_{23}^{h_{231}}}\,.$ (107)
From this we derive the formula for the action of $\Box_{x_{1}}^{k}$:
$\Box_{x_{1}}^{k}\frac{1}{x_{12}^{h_{123}}x_{13}^{h_{132}}x_{23}^{h_{231}}}=\sum^{k}_{\begin{subarray}{c}i,j=0\\ i+j\geqslant k\end{subarray}}\left[B_{i,j}C_{i,j}^{h_{123},h_{132}}\frac{x_{23}^{2(i+j-k)}}{x_{12}^{2i}x_{13}^{2j}}\right]\frac{1}{x_{12}^{h_{123}}x_{13}^{h_{132}}x_{23}^{h_{231}}}\,,$ (108)
where the coefficients are
$B_{i,j}$ $=\frac{(-1)^{i+j-k}2^{2k}k!}{(k-i)!\,(k-j)!\,(i+j-k)!}\,,$ (109)
$C_{i,j}^{a,b}$ $=\left(\frac{a}{2}\right)_{i}\left(\frac{b}{2}\right)_{j}\left(\frac{a+b}{2}-\frac{d-2}{2}+i+j-k\right)_{2k-i-j}\,.$ (110)
Since each summand on the right-hand side of (108) can be seen as a new three-point function, the formula associated with $\Box_{x_{1}}^{k}\Box_{x_{2}}^{k}\Box_{x_{3}}^{k}$ follows:
$\Box_{x_{1}}^{k}\Box_{x_{2}}^{k}\Box_{x_{3}}^{k}\frac{1}{x_{12}^{h_{123}}x_{13}^{h_{132}}x_{23}^{h_{231}}}$
=\;= $\sum^{k}_{\begin{subarray}{c}i_{1},j_{1}=0\\ i_{1}+j_{1}\geqslant k\end{subarray}}\;\;\sum^{k}_{\begin{subarray}{c}i_{2},j_{2}=0\\ i_{2}+j_{2}\geqslant k\end{subarray}}\;\;\sum^{k}_{\begin{subarray}{c}i_{3},j_{3}=0\\ i_{3}+j_{3}\geqslant k\end{subarray}}\Bigg{[}B_{i_{1},j_{1}}C_{i_{1},j_{1}}^{h_{123},h_{132}}B_{i_{2},j_{2}}C_{i_{2},j_{2}}^{h_{123}+2i_{1},h_{231}+2(k-i_{1}-j_{1})}$
×$\times B_{i_{3},j_{3}}C_{i_{3},j_{3}}^{h_{231}+2(j_{2}+k-i_{1}-j_{1}),h_{132}+2(j_{1}+k-i_{2}-j_{2})}x_{12}^{2(i_{3}+j_{3}-i_{1}-i_{2}-k)}$
×$\times x_{13}^{2(i_{2}+j_{2}-j_{1}-j_{3}-k)}x_{23}^{2(i_{1}+j_{1}-j_{2}-i_{3}-k)}\Bigg{]}\frac{1}{x_{12}^{h_{123}}x_{13}^{h_{132}}x_{23}^{h_{231}}}\,.$ (111)
Appendix B Analytic bootstrap at subleading twist
In this appendix, we check the consistency between the multiplet-recombination results and crossing symmetry at subleading twist $m=1$ using the analytic bootstrap.
Let us start with the generalized Yang-Lee edge singularity. We examine the subleading term of the crossing equation (33) in the $z\rightarrow 0$ expansion. On the left-hand side of the crossing equation, both $\mathcal{J}^{(0)}_{\ell}$ and $\mathcal{J}^{(1)}_{\ell}$ contributes to the infinite sum. The lightcone expansion of the conformal blocks are written as
$G_{\mathcal{O}}(z,\bar{z})=z^{(\Delta_{\mathcal{O}}-\ell)/2}k_{h}(\bar{z})+\sum_{i=-1}^{1}c_{i}\,z^{(\Delta_{\mathcal{O}}-\ell)/2+1}k_{h+i}(\bar{z})+O(z^{(\Delta_{\mathcal{O}}-\ell)/2+2})\,,$ (112)
For $\mathcal{O}=\mathcal{J}^{(0)}_{\ell}$, the relevant part is the subleading term in this expansion. Here $c_{i}$ are obtained by requiring that (112) solves the Casimir equation Dolan:2003hv . The results are
$c_{-1}=\frac{(d-2)\ell}{d+2\ell-4}\,,\quad c_{0}=\frac{\Delta_{\mathcal{O}}-\ell}{4}\,,\quad c_{1}=-\frac{(d-2)(\Delta_{\mathcal{O}}-1)(\Delta_{\mathcal{O}}+\ell)^{2}}{16(d-2\Delta_{\mathcal{O}}-2)(\Delta_{\mathcal{O}}+\ell-1)(\Delta_{\mathcal{O}}+\ell+1)}\,.$ (113)
We expand (112) in the anomalous dimension (41):
$G_{\mathcal{J}^{(0)}_{\ell}}(z,\bar{z})=\;$ $z^{\Delta_{\smash{\phi}}}\left(k_{h^{(0)}}(\bar{z})+\frac{1}{2}\tilde{\gamma}_{\mathcal{J}^{(0)}_{\ell}}k_{h^{(0)}}(\bar{z})\log z+O(\tilde{\gamma}_{\smash{\mathcal{J}^{(0)}_{\ell}}}^{2})\right)$
$+z^{\Delta_{\smash{\phi}}+1}\left(\sum_{i=-1}^{1}c_{i}k_{h^{(0)}+i}(\bar{z})+\frac{1}{2}\tilde{\gamma}_{\mathcal{J}^{(0)}_{\ell}}\sum_{i=-1}^{1}c_{i}k_{h^{(0)}+i}(\bar{z})\log z+O(\tilde{\gamma}_{\smash{\mathcal{J}^{(0)}_{\ell}}}^{2})\right)$
$+O(z^{(\Delta_{\smash{\mathcal{J}^{(0)}_{\ell}}}-\ell)/2+2})\,,$ (114)
where the conformal spin is $h^{(m)}\equiv h_{\smash{\mathcal{J}^{(m)}_{\ell}}}=(\Delta_{\smash{\mathcal{J}^{(m)}_{\ell}}}+\ell)/2$. For $\mathcal{O}=\mathcal{J}^{(1)}_{\ell}$, it is sufficient to consider the leading term in the $z\rightarrow 0$ expansion:
$G_{\mathcal{J}^{(1)}_{\ell}}(z,\bar{z})=z^{\Delta_{\smash{\phi}}+1}\left(k_{h^{(1)}}(\bar{z})+\frac{1}{2}\tilde{\gamma}_{\mathcal{J}^{(1)}_{\ell}}k_{h^{(1)}}(\bar{z})\log z+O(\tilde{\gamma}_{\smash{\mathcal{J}^{(1)}_{\ell}}}^{2})\right)+O(z^{(\Delta_{\smash{\mathcal{J}^{(1)}_{\ell}}}-\ell)/2+1})\,.$ (115)
At order $\epsilon^{0}$, the left-hand side of the crossing equation contains the infinite sums associated with $\mathcal{J}^{(1)}_{\ell}$ and $\mathcal{J}^{(0)}_{\ell}$:
$\sum_{\ell=0,2,4,\ldots}P_{\mathcal{J}^{(1)}_{\ell}}k_{h^{(1)}}(\bar{z})+\sum_{\ell=0,2,4,\ldots}P_{\mathcal{J}^{(0)}_{\ell}}\sum_{i=-1}^{1}c_{i}k_{h^{(0)}+i}(\bar{z})\,,$ (116)
which is present at subleading order in the $z\rightarrow 0$ expansion, i.e., at order $z^{\Delta_{\phi}+1}$. To make use of the asymptotic relation (39), we rewrite it as a single infinite sum of $\text{SL}(2,\mathbb{R})$ blcoks
$\sum_{\ell=0,2,4,\ldots}\left[P_{\mathcal{J}^{(1)}_{\ell}}+\sum_{i=-1}^{1}\left(P_{\mathcal{J}^{(0)}_{\ell}}c_{i}\right)_{\ell\rightarrow\ell-i+3}\right]k_{h^{(1)}}(\bar{z})+O(\epsilon)\,.$ (117)
The enhanced singularity on the right-hand side of the crossing equation is $\Delta_{\smash{\phi_{\text{f}}}}(\frac{1-\bar{z}}{\bar{z}})^{-\Delta_{\smash{\phi_{\text{f}}}}}$, which corresponds to the identity operator contribution. Since $\Delta_{\smash{\phi_{\text{f}}}}=2k$, the asymptotic relation implies that the coefficient in (117) is given by $2S_{-2k}(h^{(1)})$ at order $\epsilon^{0}$. Subtracting the contribution from the leading twist operators $\mathcal{J}^{(0)}_{\ell}$, we find the free squared OPE coefficients
$P_{\mathcal{J}^{(1)}_{\ell},\text{f}}=-\frac{(k-1)(2k)_{\ell+1}^{2}}{\ell!(k+\ell+1)(3k+\ell)(4k+\ell+1)_{\ell}}\,,$ (118)
where we have used the free squared OPE coefficients at leading twist (43) .
For the leading correction term in the expansion in the anomalous dimensions, the infinite sum is given by
$\sum_{\ell=0,2,4,\ldots}\left[P_{\mathcal{J}^{(1)}_{\ell}}\tilde{\gamma}_{\smash{\mathcal{J}^{(1)}_{\ell}}}+\sum_{i=-1}^{1}\left(P_{\mathcal{J}^{(0)}_{\ell}}\tilde{\gamma}_{\smash{\mathcal{J}^{(0)}_{\ell}}}c_{i}\right)_{\ell\rightarrow\ell-i+3}\right]k_{h^{(1)}}(\bar{z})\log z+O(\epsilon)\,.$ (119)
On the right-hand side of the crossing equation, the leading correction term is of order $\epsilon^{1}$ and is associated with the contribution of $\phi$. This term corresponds to the infinite sum involving anomalous dimensions on the left-hand side. For a scalar operator $\mathcal{O}$, the $\log z$ part of the conformal block takes the form Li:2020ijq
$G_{\mathcal{O}}(1-\bar{z},1-z)\big{|}_{\log z}$
$=$ $\;(1-\bar{z})^{\frac{\Delta_{\mathcal{O}}}{2}}\frac{\Gamma(\Delta_{\mathcal{O}})}{\Gamma(\Delta_{\mathcal{O}}/2)^{2}}\Big{\{}-\,_{2}F_{1}\left(\frac{\Delta_{\mathcal{O}}}{2},\frac{\Delta_{\mathcal{O}}}{2};\Delta_{\mathcal{O}}-\frac{d-2}{2};1-\bar{z}\right)$
$+\frac{\Delta_{\mathcal{O}}}{4}\left[(2-\Delta_{\mathcal{O}})\,_{2}F_{1}\left(\frac{\Delta_{\mathcal{O}}}{2},\frac{\Delta_{\mathcal{O}}}{2};\Delta_{\mathcal{O}}-\frac{d-2}{2};1-\bar{z}\right)\right.$
$+\left.\frac{(d-2)\Delta_{\mathcal{O}}}{d-2\Delta_{\mathcal{O}}-2}(1-\bar{z})\,_{2}F_{1}\left(\frac{\Delta_{\mathcal{O}}}{2}+1,\frac{\Delta_{\mathcal{O}}}{2}+1;\Delta_{\mathcal{O}}-\frac{d-4}{2};1-\bar{z}\right)\right]z+O(z^{2})\Big{\}}\,.$ (120)
The contribution from $\phi$ reads $z^{\Delta_{\phi}}(1+\Delta_{\phi}\,z+O(z^{2}))\frac{\bar{z}^{\Delta_{\smash{\phi}}}}{(1-\bar{z})^{\Delta_{\smash{\phi}}}}P_{\phi}G_{\phi}(1-\bar{z},1-z)|_{\log z\text{ part}}$. At subleading order in the lightcone expansion, this leads to the enhanced singularity given by
$P_{\phi}\frac{\Gamma(2k)}{\Gamma(k)^{2}}\Bigg{[}\left(\Delta_{\phi}+\frac{\Delta_{\phi}\left(\Delta_{\phi}-2\right)}{4}\right)\sum_{i=0}^{k-1}(-1)^{i+1}\frac{(1-2k)_{i}(k)_{i}}{i!\,(1-k)_{i}}\left(\frac{1-\bar{z}}{\bar{z}}\right)^{i-k}$
$\qquad\qquad\;-\frac{\left(d-2\right)\Delta_{\phi}^{2}}{4\left(d-2\Delta_{\phi}-2\right)}\sum_{i=0}^{k-2}(-1)^{i+1}\frac{(1-2k)_{i}(k+1)_{i}}{i!\,(2-k)_{i}}\left(\frac{1-\bar{z}}{\bar{z}}\right)^{i-k+1}$
[…]
where we have only kept the terms of order $\epsilon^{1}$. Using (24), (37) and (39), we find the coefficients in the infinite sum (119). The anomalous dimensions $\tilde{\gamma}_{\smash{\mathcal{J}^{(1)}_{\ell}}}$ are obtained by subtracting the contribution from $\mathcal{J}^{(0)}_{\ell}$. The results agree with those from the multiplet recombination.
For the generalized Potts model, the free squared OPE coefficients are related to those at $N=1$ by
$P_{\mathcal{J}^{(1)}_{\ell},\text{f}}^{\text{ generic }N}=P_{\mathcal{J}^{(1)}_{\ell},\text{f}}^{\;N=1}\times\begin{cases}\;\frac{1}{N}\qquad&\text{trivial}\\ \;\frac{1}{(N-1)(N+1)^{2}}&\text{standard}\\ \;1&\text{symmetric traceless}\\ \;-1&\text{antisymmetric}\,,\end{cases}$ (122)
which is the $m=1$ version of (78) at leading twist. For the anomalous dimensions, we consider the enhanced singularity given by (B). From this we extract the coefficients in the infinite sum of the form (119). Here we need to account for the $N$-dependent factors in the crossing equations (D)-(D). As a result, the coefficients in the infinite sums have the following relation with those in the Yang-Lee case:
[…]
We subtract the contribution from $\mathcal{J}^{(0)}_{\ell}$, and obtain the anomalous dimensions in the Potts model. The results are again in agreement with those from the multiplet recombination.
Finally, let us consider the generalized $\phi^{2n+1}$ theory with $n>1$. At order $\alpha^{0}$, the enhanced singularity at subleading order in the lightcone expansion is given by $\Delta_{\smash{\phi_{\text{f}}}}(\frac{1-\bar{z}}{\bar{z}})^{-\Delta_{\smash{\phi_{\text{f}}}}}$, with the scaling dimension $\Delta_{\phi_{\text{f}}}=\frac{2k}{2n-1}$. We obtain the free squared OPE coefficient
$P_{\smash{\mathcal{J}_{\ell}^{(1)}},\text{f}}=\frac{(k-1)(2n-1)^{2}\left(\frac{2k}{2n-1}\right)_{\ell+1}^{2}}{\ell!(k-\ell+2n(k+\ell))(\ell-3k+2n(k-\ell-1)+1)\left(\frac{4k}{2n-1}+\ell+1\right)_{\ell}}\,.$ (124)
At order $\alpha^{2}$, we consider the inversion formula, which is associated with the double discontinuity given by
$z^{\Delta_{\phi^{2n-1}}}(1+\Delta_{\phi^{2n-1}}\,z+O(z^{2}))\frac{\bar{z}^{\Delta_{\phi^{2n-1}}}}{(1-\bar{z})^{\Delta_{\phi^{2n-1}}}}P_{\phi^{2n-1}}G_{\phi^{2n-1}}(1-\bar{z},1-z)\Big{|}_{\log z\text{ part}}\,.$ (125)
Multiplying by the factor $2\sin^{2}(\pi k\,\frac{2n-3}{2n-1})$, we obtain the corresponding dDsic, as in (4.2). At subleading order in the $z\rightarrow 0$ expansion, we find the anomalous dimensions of $\mathcal{J}^{(1)}_{\ell}$ using the leading-twist anomalous dimension, as well as the free squared OPE coefficients (100) and (124). The results are consistent with those obtained from the multiplet recombination.
In conclusion, we confirm the consistency between the multiplet-recombination results and crossing symmetry at subleading twist.
Appendix C Representation theory of $S_{N+1}$
The ($N+1$)-component vectors transform naturally under $\tilde{P}^{\alpha\beta}\in S_{N+1}$, where $\tilde{P}^{\alpha\beta}$ is a permutation matrix. However, this does not lead to an irreducible representation, since the ($N+1$)-component vector $(1,1,\ldots,1)$ forms the basis for a 1D invariant subspace. To construct a nontrivial irreducible representation, it is convenient to think of $S_{N+1}$ as the symmetry group of an $N$-simplex, which has $N+1$ vertices. The vertices are in one-to-one correspondence with the vectors $e_{a}^{\alpha}$ which connect the center and the vertices. The Latin index $a$ ranges from 1 to $N$, as the simplex is embedded in an $N$-dimensional space. The Greek index $\alpha$ ranges from 1 to $N+1$ because there are $N+1$ vertices. The vectors satisfy the relations Zia:1975ha  [^28] The normalization in Zia:1975ha is different from ours: $e_{a,\text{ours}}^{\alpha}=\sqrt{N}\;e_{a,\text{theirs}}^{\alpha}$.
$\sum_{\alpha=1}^{N+1}e_{a}^{\alpha}=0\,,\qquad\sum_{\alpha=1}^{N+1}e_{a}^{\alpha}e_{b}^{\alpha}=(N+1)\delta_{ab}\,,\qquad\sum_{a=1}^{N}e_{a}^{\alpha}e_{a}^{\beta}=(N+1)\delta^{\alpha\beta}-1\,.$ (126)
An $N$-component vector $v_{a}$ can be projected onto the vectors $e_{a}^{1},e_{a}^{2},\ldots,e_{a}^{N+1}$ (see figure 14), and the results can be viewed as an $(N+1)$-component vector which transforms under $\tilde{P}^{\alpha\beta}$. Then, we write the vector after the transformation as an $N$-component vector again. In short, the transformation rule for $N$-component vectors is given by
$v_{a}=\sum_{b=1}^{N}P_{ab}\,v_{b}\,,\qquad P_{ab}=\frac{1}{N+1}\sum_{\alpha,\beta=1}^{N+1}e^{\alpha}_{a}\tilde{P}^{\alpha\beta}e^{\beta}_{b}\,.$ (127)
In this way, the space of the $N$-component vectors furnishes an irreducible representation of $S_{N+1}$, which is isomorphic to the standard representation of $S_{N+1}$. [^29] The standard representation of $S_{N+1}$ consists of ($N+1$)-component vectors $u^{\alpha}$. They subject to the condition that the sum of the components vanishes $\sum_{\alpha=1}^{N+1}u^{\alpha}=0$. We do not distinguish between the two isomorphic vector spaces, and refer to the space of $N$-component vectors as the standard representation.
$e^{1}_{a}$$e^{2}_{a}$$e^{3}_{a}$
Figure 14: Projection of a two-component vector. Using the vectors $e^{\alpha}_{a}$ in the 2-simplex (equilateral triangle), the 2D vector in red is associated with three components.
The dimensions of the corresponding representations $V_{(\ldots)}$ are given by the hook length formula $\frac{(N+1)!}{\prod(\text{hook lengths})}$. We have
$\dim V_{(N+1)}$ $=\frac{(N+1)!}{(N+1)\times N\times\ldots\times 1}=1\,,$ (128)
$\dim V_{(N,1)}$ $=\frac{(N+1)!}{(N+1)\times(N-1)\times(N-2)\times\ldots\times 1\times 1}=N\,,$ (129)
$\dim V_{(N,1)}$ $=\frac{(N+1)!}{N\times(N-1)\times(N-3)\times\ldots\times 1\times 2\times 1}=\frac{(N+1)(N-2)}{2}\,,$ (130)
$\dim V_{(N-1,1^{2})}$ $=\frac{(N+1)!}{(N+1)\times(N-2)\times(N-3)\times\ldots\times 1\times 2\times 1}=\frac{N(N-1)}{2}\,.$ (131)
More concretely, using the projectors in appendix D, we obtain some examples of the bilinear operators:
$\sum_{c,d=1}^{N}t^{(1)}_{abcd}\,\phi_{c}\phi_{d}$ $\rightarrow\sum_{a=1}^{N}\phi_{a}\phi_{a}=\phi^{2}\,,$ (132)
$\sum_{c,d=1}^{N}t^{(2)}_{abcd}\,\phi_{c}\phi_{d}$ $\rightarrow\sum_{b,c=1}^{N}T_{abc}\,\phi_{b}\phi_{c}\,,$ (133)
$\sum_{c,d=1}^{N}t^{(3)}_{abcd}\,\phi_{c}\phi_{d}$ $\rightarrow(N+1)(N-1)\phi_{a}\phi_{b}+\frac{N+1}{N}\delta_{ab}\phi^{2}-\sum_{\alpha=1}^{N+1}\sum_{c,d=1}^{N}e_{a}^{\alpha}e_{b}^{\alpha}e_{c}^{\alpha}e_{d}^{\alpha}\phi_{c}\phi_{d}\,,$ (134)
$\sum_{c,d=1}^{N}t^{(4)}_{abcd}\,\phi_{c}\partial_{\mu}\phi_{d}$ $\rightarrow\phi_{a}\partial_{\mu}\phi_{b}-\phi_{b}\partial_{\mu}\phi_{a}\,.$ (135)
The last two lines follow directly from the definition of $t_{abcd}^{(i)}$. The first line is obtained by removing the redundancy $\delta_{ab}$. The second line follows from
$\sum_{c,d=1}^{N}t^{(2)}_{abcd}\phi_{c}\phi_{d}=(N+1)\left(\sum_{\alpha=1}^{N+1}\sum_{c,d=1}^{N}e_{a}^{\alpha}e_{b}^{\alpha}e_{c}^{\alpha}e_{d}^{\alpha}\phi_{c}\phi_{d}-(N+1)\delta_{ab}\phi^{2}\right)=\sum_{c,i,j=1}^{N}T_{abc}T_{cij}\phi_{i}\phi_{j}\,,$ (136)
which implies that $\sum_{c,d=1}^{N}t^{(2)}_{abcd}\phi_{c}\phi_{d}$ is a redundant description of $\sum_{b,c=1}^{N}T_{abc}\phi_{b}\phi_{c}$. We thus focus on the latter one.
Comparing the dimensions with those of (128), (129), (130) and (131), we conclude that (132), (133), (134) and (135) correspond to the Young diagrams $(N+1)$, $(N,1)$, $(N-1,2)$ and $(N-1,1^{2})$. [^30] The dimension of (134) follows from the symmetric tracelessness and the constraints $\sum_{a,b=1}^{N}T_{iab}\sum_{c,d=1}^{N}t^{(3)}_{abcd}\phi_{c}\phi_{d}=0\,,$ (137) where $i=1,2,\ldots,N$. From these examples, we see that the tensor product decomposition resembles that in the O($N$) model, except that here the symmetric traceless part further decomposes into two irreducible representations $V_{(N,1)}$ and $V_{(N-1,2)}$. Since $V_{(N,1)}$ is the standard representation, we use the term “symmetric traceless representation” to refer to $V_{(N-1,2)}$ below. Moreover, we refer to $V_{(N+1)}$ and $V_{(N-1,1^{2})}$ as the “trivial representation” and the “antisymmetric representation”. [^31] The antisymmetric representation $V_{(N-1,1^{2})}$ is the exterior power $\Lambda^{2}V_{(N,1)}$ of the standard representation $V_{(N,1)}$.
Appendix D Tensor structures
The tensor product decomposition of $(N,1)^{4}$ contains the trivial representation four times, so there are four linearly independent rank-4 invariant tensors wallace1978spin . Using $e_{a}^{\alpha}$ defined by (126), we identify four rank-4 $S_{N+1}$-invariant tensors: $\delta_{ab}\delta_{cd}$, $\delta_{ac}\delta_{bd}$, $\delta_{ad}\delta_{bc}$ and $\sum_{\alpha=1}^{N+1}e_{a}^{\alpha}e_{b}^{\alpha}e_{c}^{\alpha}e_{d}^{\alpha}$. We then use these tensors to define a basis for the four-point function $\langle\phi_{a}(x_{1})\phi_{b}(x_{2})\phi_{c}(x_{3})\phi_{d}(x_{4})\rangle$:
$t_{abcd}^{(1)}$ $=\delta_{ab}\delta_{cd}\,,$ (138)
$t_{abcd}^{(2)}$ $=(N+1)\left(\sum_{\alpha=1}^{N+1}e_{a}^{\alpha}e_{b}^{\alpha}e_{c}^{\alpha}e_{d}^{\alpha}-(N+1)\delta_{ab}\delta_{cd}\right)\,,$ (139)
$t_{abcd}^{(3)}$ $=\frac{1}{2}\left(\delta_{ac}\delta_{bd}+\delta_{ad}\delta_{bc}\right)+\frac{1}{N(N-1)}\delta_{ab}\delta_{cd}-\frac{1}{(N+1)(N-1)}\sum_{\alpha=1}^{N+1}e_{a}^{\alpha}e_{b}^{\alpha}e_{c}^{\alpha}e_{d}^{\alpha}\,,$ (140)
$t_{abcd}^{(4)}$ $=\frac{1}{2}\left(\delta_{ac}\delta_{bd}-\delta_{ad}\delta_{bc}\right)\,.$ (141)
For $i=1,2,3,4$, the tensor structures $t_{abcd}^{(i)}$ are associated with the intermediate operators in the trivial, standard, symmetric tracless and antisymmetric representations. The four tensors can also be used as projectors, which satisfy
$\sum_{c,d=1}^{N}t_{a_{1}b_{1}c\,d}^{(i)}\,t_{c\,d\,a_{2}b_{2}}^{(j)}\propto\delta^{ij}t_{a_{1}b_{1}a_{2}b_{2}}^{(i)}\,.$ (142)
The normalization of $t_{abcd}^{(i)}$ is chosen as follows. The free correlator $\langle\phi_{a,\text{f}}(x_{1})\phi_{b,\text{f}}(x_{2})\phi_{c,\text{f}}(x_{3})$ $\phi_{d,\text{f}}(x_{4})\rangle$ is computed using Wick’s theorem, yielding four parts that correspond to the four tensor structures. We require that the conformal block decomposition of each part is precisely (34), with the same normalization of OPE coefficients and conformal blocks. We use these projectors to construct bilinear operators in the irreducible representations, as in (132), (133), (134) and (135).
Using the tensor structures defined above, the four-point function is written as
$\langle\phi_{a}(x_{1})\phi_{b}(x_{2})\phi_{c}(x_{3})\phi_{d}(x_{4})\rangle=\frac{1}{x_{12}^{2\Delta_{\smash{\phi}}}x_{34}^{2\Delta_{\smash{\phi}}}}\sum_{i=1}^{4}g^{(i)}(z,\bar{z})\,t_{abcd}^{(i)}\,.$ (143)
The crossing equations read
$g^{(1)}(z,\bar{z})=\;$ $\frac{z^{\Delta_{\smash{\phi}}}\bar{z}^{\Delta_{\smash{\phi}}}}{(1-z)^{\Delta_{\smash{\phi}}}(1-\bar{z})^{\Delta_{\smash{\phi}}}}\Bigg{(}\frac{g^{(1)}(1-\bar{z},1-z)}{N}+\frac{(N-1)(N+1)^{2}}{N}g^{(2)}(1-\bar{z},1-z)$
$+\frac{(N-2)(N+1)}{2N^{2}}g^{(3)}(1-\bar{z},1-z)+\frac{1-N}{2N}g^{(4)}(1-\bar{z},1-z)\Bigg{)}\,,$ (144)
$g^{(2)}(z,\bar{z})=\;$ $\frac{z^{\Delta_{\smash{\phi}}}\bar{z}^{\Delta_{\smash{\phi}}}}{(1-z)^{\Delta_{\smash{\phi}}}(1-\bar{z})^{\Delta_{\smash{\phi}}}}\Bigg{(}\frac{g^{(1)}(1-\bar{z},1-z)}{(N-1)(N+1)^{2}}+\frac{N-2}{N-1}g^{(2)}(1-\bar{z},1-z)$
$-\frac{N-2}{2N(N+1)(N-1)^{2}}g^{(3)}(1-\bar{z},1-z)+\frac{g^{(4)}(1-\bar{z},1-z)}{2(N-1)(N+1)^{2}}\Bigg{)}\,,$ (145)
$g^{(3)}(z,\bar{z})=\;$ $\frac{z^{\Delta_{\smash{\phi}}}\bar{z}^{\Delta_{\smash{\phi}}}}{(1-z)^{\Delta_{\smash{\phi}}}(1-\bar{z})^{\Delta_{\smash{\phi}}}}\Bigg{(}g^{(1)}(1-\bar{z},1-z)-(N+1)^{2}g^{(2)}(1-\bar{z},1-z)$
$+\frac{N^{2}-N+2}{2N(N-1)}g^{(3)}(1-\bar{z},1-z)+\frac{g^{(4)}(1-\bar{z},1-z)}{2}\Bigg{)}\,,$ (146)
$g^{(4)}(z,\bar{z})=\;$ $\frac{z^{\Delta_{\smash{\phi}}}\bar{z}^{\Delta_{\smash{\phi}}}}{(1-z)^{\Delta_{\smash{\phi}}}(1-\bar{z})^{\Delta_{\smash{\phi}}}}\Bigg{(}-g^{(1)}(1-\bar{z},1-z)+(N+1)^{2}g^{(2)}(1-\bar{z},1-z)$
$+\frac{(N-2)(N+1)}{2N(N-1)}g^{(3)}(1-\bar{z},1-z)+\frac{g^{(4)}(1-\bar{z},1-z)}{2}\Bigg{)}\,.$ (147)
The analytic bootstrap in subsection 3.2 is based on these crossing equations.
Appendix E Ratios of OPE coefficients
In the matching conditions (2.1) and (4.1), we encounter the ratios of OPE coefficients $\frac{\lambda_{\phi^{2n}\phi^{2n}\mathcal{J},\text{f}}}{\lambda_{\phi\phi\mathcal{J},\text{f}}}$, where $n\geqslant 1$. Moreover, the ratio $\frac{\lambda_{KK\mathcal{J},\text{f}}}{\lambda_{1,1,\mathcal{J},\text{f}}}$ in (68) depends on $N$, the number of components of $\phi_{a}$. In this appendix, we determine these ratios by inspecting the conformal block decompositions of four-point functions.
Let us first focus on the $N=1$ case. We only need the free theory data here, and we can consider the ratios with $n=1$ and $n>1$ together. The four-point function of identical scalars reads
$\langle\phi_{\text{f}}(x_{1})\phi_{\text{f}}(x_{2})\phi_{\text{f}}(x_{3})\phi_{\text{f}}(x_{4})\rangle=\left[1+\left(z\bar{z}\right)^{\Delta_{\phi_{\text{f}}}}+\left(\frac{z\bar{z}}{(1-z)(1-\bar{z})}\right)^{\Delta_{\phi_{\text{f}}}}\right]\mathcal{K}(\Delta_{i},x_{i})\,.$ (148)
The mixed correlator of interest is
$\langle\phi_{\text{f}}^{2n}(x_{1})\phi_{\text{f}}^{2n}(x_{2})\phi_{\text{f}}(x_{3})\phi_{\text{f}}(x_{4})\rangle$
=\;= $\left[(2n)!+(2n)(2n)!\left(\left(z\bar{z}\right)^{\Delta_{\phi_{\text{f}}}}+\left(\frac{z\bar{z}}{(1-z)(1-\bar{z})}\right)^{\Delta_{\phi_{\text{f}}}}\right)\right]\mathcal{K}(\Delta_{i},x_{i})\,,$ (149)
where the factors $(2n)!$ and $(2n)(2n)!$ come from Wick contraction. The kinematic factors are given by
$\mathcal{K}(\Delta_{i},x_{i})=\frac{1}{x_{12}^{\Delta_{1}+\Delta_{2}}x_{34}^{\Delta_{3}+\Delta_{4}}}\left(\frac{x_{24}}{x_{14}}\right)^{\Delta_{1}-\Delta_{2}}\left(\frac{x_{14}}{x_{13}}\right)^{\Delta_{3}-\Delta_{4}}\,,$ (150)
where $\Delta_{i}$ is the scaling dimension of the operator at position $x_{i}$. In (148) and (E), the terms in the square brackets have the same functional form in $z$ and $\bar{z}$, so their conformal block decompositions differ by a factor $(2n)(2n)!$. In other words, there exists a relation between the OPE coefficients
$(2n)(2n)!\lambda_{\phi\phi\mathcal{J}}^{2}=\lambda_{\phi^{2n}\phi^{2n}\mathcal{J}}\lambda_{\phi\phi\mathcal{J}}\,,$ (151)
which yields the ratio
$\frac{\lambda_{\phi^{2n}\phi^{2n}\mathcal{J},\text{f}}}{\lambda_{\phi\phi\mathcal{J},\text{f}}}=(2n)(2n)!\,.$ (152)
In particular, we have $\frac{\lambda_{\phi^{2n}\phi^{2n}\mathcal{J},\text{f}}}{\lambda_{\phi\phi\mathcal{J},\text{f}}}=4$ at $n=1$.
Now we consider the case of $N$-component fields. The relevant four-point functions are
$\langle\phi_{a,\text{f}}(x_{1})\phi_{b,\text{f}}(x_{2})\phi_{c,\text{f}}(x_{3})\phi_{d,\text{f}}(x_{4})\rangle$
=\;= $\Bigg{[}\left(1+\frac{\left(z\bar{z}\right)^{\Delta_{\phi_{\text{f}}}}\left(1+\left((1-z)(1-\bar{z})\right)^{-\Delta_{\phi_{\text{f}}}}\right)}{N}\right)t_{abcd}^{(1)}$
$\quad+\frac{\left(z\bar{z}\right)^{\Delta_{\phi_{\text{f}}}}\left(1+\left((1-z)(1-\bar{z})\right)^{-\Delta_{\phi_{\text{f}}}}\right)}{(N-1)(N+1)^{2}}t_{abcd}^{(2)}$
$\quad+\left(z\bar{z}\right)^{\Delta_{\phi_{\text{f}}}}\left(1+\left((1-z)(1-\bar{z})\right)^{-\Delta_{\phi_{\text{f}}}}\right)t_{abcd}^{(3)}$
$\quad+\left(z\bar{z}\right)^{\Delta_{\phi_{\text{f}}}}\left(1-\left((1-z)(1-\bar{z})\right)^{-\Delta_{\phi_{\text{f}}}}\right)t_{abcd}^{(4)}\Bigg{]}\mathcal{K}(\Delta_{i},x_{i})$ (153)
and
$\langle K_{a,\text{f}}(x_{1})K_{b,\text{f}}(x_{2})\phi_{c,\text{f}}(x_{3})\phi_{d,\text{f}}(x_{4})\rangle$
=\;= $\Bigg{[}\left(2(N-1)(N+1)^{2}+\frac{4(N-1)(N+1)^{2}\left(z\bar{z}\right)^{\Delta_{\phi_{\text{f}}}}\left(1+\left((1-z)(1-\bar{z})\right)^{-\Delta_{\phi_{\text{f}}}}\right)}{N}\right)t_{abcd}^{(1)}$
$\quad+\frac{4(N-2)\left(z\bar{z}\right)^{\Delta_{\phi_{\text{f}}}}\left(1+\left((1-z)(1-\bar{z})\right)^{-\Delta_{\phi_{\text{f}}}}\right)}{N-1}t_{abcd}^{(2)}$
$\quad-4(N+1)^{2}\left(z\bar{z}\right)^{\Delta_{\phi_{\text{f}}}}\left(1+\left((1-z)(1-\bar{z})\right)^{-\Delta_{\phi_{\text{f}}}}\right)t_{abcd}^{(3)}$
$\quad-4(N+1)^{2}\left(z\bar{z}\right)^{\Delta_{\phi_{\text{f}}}}\left(1-\left((1-z)(1-\bar{z})\right)^{-\Delta_{\phi_{\text{f}}}}\right)t_{abcd}^{(4)}\;\Bigg{]}\mathcal{K}(\Delta_{i},x_{i})\,.$ (154)
[…]
$R=\frac{\lambda_{KK\mathcal{J},\text{f}}}{\lambda_{\phi\phi\mathcal{J},\text{f}}}=\begin{cases}\;4(N-1)(N+1)^{2}\qquad&\text{trivial}\\ \;4(N-2)(N+1)^{2}&\text{standard}\\ \;-4(N+1)^{2}&\text{symmetric traceless}\\ \;-4(N+1)^{2}&\text{antisymmetric}\,.\end{cases}$ (155)
Appendix F An OPE coefficient in generalized $\phi^{2n+1}$ theory ($n>1$)
We are interested in the OPE coefficient $\lambda_{\phi\phi\phi^{2n-1}}$ in $\phi\times\phi\sim\phi^{2n-1}$. This coefficient is useful in the analytic bootstrap, as in (4.2). We consider the matching condition
$\lim_{\epsilon\rightarrow 0}\left(\alpha^{-2}\langle\Box^{k}\phi(x_{1})\phi(x_{2})\phi^{2n-1}(x_{3})\rangle\right)=\langle\phi^{2n}_{\text{f}}(x_{1})\phi_{\text{f}}(x_{2})\phi^{2n-1}(x_{3})\rangle\,.$ (156)
At the leading order in the $x_{3}\rightarrow\infty$ expansion, the condition above yields
$\lim_{\epsilon\rightarrow 0}$ $\Bigg{[}\alpha^{-2}2^{2k}\left(\Delta_{\smash{\phi}}-\frac{\Delta_{\phi^{2n-1}}}{2}\right)_{k}\left(\Delta_{\smash{\phi}}-\frac{\Delta_{\phi^{2n-1}}}{2}-\frac{d-2}{2}\right)_{k}\frac{\lambda_{\phi\phi\phi^{2n-1}}}{x_{12}^{2\Delta_{\smash{\phi}}+2k-\Delta_{\smash{\phi^{2n-1}}}}|x_{3}|^{2\Delta_{\smash{\phi^{2n-1}}}}}$
$\;\;+O(\,|x_{3}|^{-2\Delta_{\smash{\phi^{2n-1}}}-1})\Bigg{]}=\frac{\lambda_{\phi^{2n}\phi\phi^{2n-1},\text{f}}}{x_{12}^{2\Delta_{\smash{\phi^{2n}_{\text{f}}}}-\Delta_{\smash{\phi^{2n-1}_{\text{f}}}}}|x_{3}|^{2\Delta_{\smash{\phi^{2n-1}_{\text{f}}}}}}+O(\,|x_{3}|^{-2\Delta_{\smash{\phi^{2n-1}_{\text{f}}}}-1})\,.$ (157)
The constraint on $\lambda_{\phi\phi\phi^{2n-1}}$ reads
$2^{2k}\left(\frac{(3-2n)\left(d_{\text{u}}-2k\right)}{4}\right)_{k}\left(\frac{(3-2n)\left(d_{\text{u}}-2k\right)}{4}-\frac{d_{\text{u}}-2}{2}\right)_{k}\lim_{\epsilon\rightarrow 0}\left(\alpha^{-1}\lambda_{1,1,2n-1}\right)=(2n)!\,,$ (158)
where the upper critical dimension is given in (81). The solution is
$\lambda_{\phi\phi\phi^{2n-1}}=\frac{(2n)!\,\alpha}{2^{2k}(1-2k)_{k}\left(k\frac{3-2n}{2n-1}\right)_{k}}+O(\alpha^{2})\,.$ (159)
According to the convention (37) in the analytic bootstrap, we obtain
$P_{\phi^{2n-1}}=\frac{\lambda_{\phi\phi\phi^{2n-1}}^{2}}{(2n-1)!}\,.$ (160)
We have used the expressions (159) and (160) in (4.2).
Appendix G Generalized self-avoiding walks and loop-erased random walks
As prominent models of random walks, we consider the self-avoiding walks and loop-erased random walks that play important roles in physics. Interestingly, they can be associated with the O($N$) symmetric $\phi^{4}$ theory. In the early 1970s, it was discovered that self-avoiding walks correspond to the $N\rightarrow 0$ limit deGennes:1972zz . Later, it was shown that loop-erased random walks are related to the $N\rightarrow-2$ limit duplantier1992loop ; Wiese:2018dow ; helmuth2020loop . In this appendix, we discuss these limits using the results from our previous paper Guo:2023qtt .
In Guo:2023qtt , we studied the generalized $\phi^{2n}$ theory with O($N$) global symmetry.  [^32] It may be interesting to consider the $n>2$ counterpart of the special $N$ limits. The anomalous dimensions of $\phi$ at $n=2$ reads
$\gamma_{\phi}=\frac{(-1)^{k+1}(N+2)(k+1)_{k}}{8(N+8)^{2}k^{2}(2k+1)_{k-1}}\epsilon^{2}+O(\epsilon^{3})\,,$ (161)
In the limits $N\rightarrow 0$ and $N\rightarrow-2$, the formula reduces to
$\gamma_{\phi}\big{|}_{N\rightarrow 0}$ $=\frac{(-1)^{k+1}(k+1)_{k}}{256k^{2}(2k+1)_{k-1}}\epsilon^{2}+O(\epsilon^{3})\,,$ (162)
$\gamma_{\phi}\big{|}_{N\rightarrow-2}$ $=O(\epsilon^{3})\,.$ (163)
Moreover, the anomalous dimension of the symmetric traceless bilinear operator $\phi_{\text{T}}^{2}$ is
$\gamma_{\phi_{\text{T}}^{2}}=\frac{2}{N+8}\epsilon+O(\epsilon^{2})\,,$ (164)
which is independent of $k$ at leading order in the $\epsilon$ expansion. We derive the fractal dimension (or Hausdorff dimension) Shimada:2015gda
$D_{\text{F}}=d-\Delta_{\phi_{\text{T}}^{2}}=2-\frac{2}{N+8}\epsilon+O(\epsilon^{2})\,,$ (165)
where spacetime dimension is $d=4k-\epsilon$ and the fractal dimension $D_{\text{F}}$ is $k$-independent to order $\epsilon$. In particular, the limits $N\rightarrow 0$ and $N\rightarrow-2$ yield
$D_{\text{F,self-avoiding}}=2-\frac{\epsilon}{4}+O(\epsilon^{2})\,,\qquad D_{\text{F,loop-erased}}=2-\frac{\epsilon}{3}+O(\epsilon^{2})\,.$ (166)